### v1

# ZeroMerge: Parameter-Free KV Cache Compression for Memory-Efficient Long-Context LLMs

Abstract
The linear growth of key-value (KV) cache memory and quadratic computational complexity pose significant bottlenecks for large language models (LLMs) in long-context processing. While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often suffer from irreversible information loss or require costly parameter retraining. We propose ZeroMerge, a dynamic zero-shot compression framework that achieves efficient cache management through three key innovations: (1) Fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) A residual merging mechanism that preserves critical context through compensated attention scoring, and (3) Parameter-free adaptation compatible with diverse LLM architectures without retraining. Comprehensive evaluations across LLaMA-2 model demonstrate that ZeroMerge maintains full-cache performance at 5% compression ratios while doubling inference throughput at 40K token lengths. The method effectively balances memory efficiency, generation quality, and deployment flexibility, advancing practical long-context LLM applications. The code is available at https://github.com/SusCom-Lab/ZeroMerge.
Machine Learning, ICML
1 Introduction
The advancement of the Transformer architecture has brought about revolutionary changes in handling sequence data, with large language models (LLMs) being a prime illustration. LLMs have achieved significant progress, surpassing humans in diverse applications. However, handling long contexts has posed challenges, particularly the linear expansion of the KV cache in proportion to the sequence length. For instance, the deployment of a Llama 2 7B model consumes 26GB of memory. When the token length reaches 32K, the 32G of memory occupied by the KV cache becomes the main memory consumer. This, along with the quadratic computational complexity, has impeded LLM integration and performance, spurring research on optimizing the KV cache for enhanced inference efficiency.
Abundant innovative studies on architecture compression alleviate the shortage of memory (Sarah et al., 2024; Ashkboos et al., 2024). Parameter quantization techniques(Dettmers et al., 2022; Xiao et al., 2023a) further reduce memory footprint by representing KV cache with low-bit precision (e.g., 4-bit integers), though at the cost of numerical accuracy preservation. More efforts focus on the management of the KV cache to directly address this challenge. Some sparse methods(Zhang et al., 2024b, a) reserve partial top-important tokens and eliminate others to mitigate the harm to the original generation quality. These selection criteria are based on the historical contribution in the decoding procedure, spatio-temporal relationship, or data distribution. One-sided and short-sighted pre-pruning strategies inevitably disrupt the distribution of KV embeddings and introduce unrecoverable information loss. Hybrid approaches combining quantization and sparsity (Dettmers et al., 2023) attempt to balance accuracy and compression ratio, yet still require delicate bit allocation strategies. Low rank compression techniques are utilized to correct the systematic drawback in some works (Dong et al., 2024). However, the newly introduced architectures or the retraining requirement of the model parameters constitutes a significant burden. One desirable KV management strategy should 1) strictly control the memory consumption to break the context length limits 2) minimize the generation quality degradation 3) be easily applied to various LLM backbones without extra refinement cost.
Motivated by these goals, we proposed ZeroMerge, a more efficient and comprehensive dynamic zero-shot KV management algorithm.
  1. 1.
    Fine-grained memory allocation. ZeroMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and its own data distribution characteristics to conduct fine-grained memory management at the head level, making the method more adaptable to different types of long text tasks. It has achieved superior generation quality compared to the sparse strategy and is comparable to the retrained model on models. The performance of 5% of the KV cache bugdet is still on par with the complete model.
  2. 2.
    Low-cost adaptation and deployment. ZeroMerge does not introduce any model parameters and does not require retraining or model fine-tuning to adapt to different compression ratios or various task types. It is adapted to various mainstream model architectures such as Llama, Falcon, and Mistral.
  3. 3.
    Efficient compression strategy. ZeroMerge adopts a light weight operational overhead of linear operational complexity and achieves more than twice the throughput of the original model at a token length of 50K. We have proved that the cache compression mechanism adopted by ZeroMerge does not drown out the original contribution of those intact tokens as the processing sequence grows.
2 Related work
Figure 1: Comparative analysis of KV cache optimization strategies: (a) Model architecture compression reduces both model parameters and KV cache size via multi-query attention heads; (b) Parameter/data quantization trades numerical precision for cache footprint reduction through low-bit representations; (c) Sparse approximation directly addresses memory-compute growth by evicting low-contribution tokens, albeit inducing attention distribution drift; (d) Token merge mitigates sparse caching errors by learning residuals between full and sparse attention outputs, yet introduces overhead from auxiliary networks or fine-tuning; (e) ZeroMerge (proposed) achieves zero-shot compression competitive with sparse+residual methods without added parameters, while demonstrating superior generalization across task domains and compression ratios.
The phenomenon of emergence in large-scale models has ignited new enthusiasm and provided fresh directions for exploring the nonlinear leaps in reasoning and generative capabilities of LLMs—driven by the expansion of model scale and complexity. The unprecedented parameter sizes and intensive computational demands have endowed LLMs with remarkable reasoning and generative abilities, yet they also pose significant challenges, hindering their cost-effective and large-scale deployment across various scenarios. Numerous efforts have been made to address the issue.
2.1 Model architecture compression
Works such as MiniLLM (Gu et al., 2024) and SliceGPT (Ashkboos et al., 2024) have been specifically devoted to ensure that the performance of sample tasks remains intact for more compact models. MiniLLM focuses on knowledge distillation of large language models, aiming to transfer the knowledge from large and powerful models to smaller ones, thereby reducing the number of parameters in the latter. On the other hand, SliceGPT adopts a different approach by replacing weight matrix with a smaller dense matrix. Through these innovative methods, they both strive to strike a balance between controlling the parameter count and maintaining satisfactory performance on relevant tasks. In the research of attention mechanisms, (Voita et al., 2019) pointed out that the contributions of different attention heads to the overall performance of the model in the multi-head self-attention mechanism are not equal. It also provided a novel pruning method at the attention head level. Similarly, a large number of works related to Multi-Query Attention (MQA) enable different attention “head” to share a set of “keys” and “values”, thus significantly reducing the size of these tensors and the memory bandwidth requirements for incremental decoding (Shazeer, 2019; Ainslie et al., 2023).
2.2 Parameter and data quantization
Quantization techniques have emerged as a crucial approach to address the computational and storage challenges of large language models. GPTQ(Frantar et al., 2022) is a one-shot weight quantization method based on approximate second-order information. It can quantize large GPT models with a significant number of parameters in a relatively short time and with minimal accuracy degradation. This enables models to be executed on a single GPU for generative inference and provides end-to-end inference speedups when using different GPUs. INT-FlashAttention(Chen et al., 2024) is the first INT8 quantization architecture compatible with the forward workflow of FlashAttention. It significantly improves the inference speed of FlashAttention on Ampere GPUs by implementing fully INT8 activations and general matrix-multiplication kernels. It also serves as a general token-level post-training quantization framework compatible with other data formats. They not only reduce the model size but also enhance the inference speed and manage quantization errors, making large language models more accessible and efficient for various applications.
2.3 Sparse approximation
Other studies directly focus on the KV cache, another performance bottleneck induced by long sequence tasks, striving to fundamentally address the issues of storage growth and quadratic computational complexity that accompany the increase in sequence sample length. The universality of sparse attention values (Ji et al., 2021; Deng et al., 2024) has provided a new direction for improving the computational efficiency and reducing the memory footprint of models. (Roy et al., 2021) dynamically routes the input tokens to different attention heads based on their content, enabling each attention head to focus only on the relevant parts, thus enhancing the computational efficiency while maintaining the model’s performance. (Wang et al., 2020) takes advantage of the low rank property and projects it into a low-order matrix, thereby avoiding the attention layer with quadratic computational complexity. Another token aware sparse approach, also known as eviction strategy, employ diverse token importance metrics to address memory bottlenecks. The foundational work by (Xiao et al., 2023b) identifies the attention sink phenomenon, where initial tokens disproportionately attract attention scores regardless of positional distance. Building on this insight, StreamingLLM combines a sliding temporal window with persistent retention of initial token KV states, enabling stable infinite-length generation. Similarly, LM-infinite (Han et al., 2024) introduces Lambda-shaped attention masks to differentially strengthen local (recent tokens) and global (starting tokens) semantic modeling. H2O (Zhang et al., 2024b)proposes Heavy Hitter token detection through cumulative attention score thresholds, dynamically evicting non-critical tokens during decoding. Simultaneously, PyramidKV(Zhang et al., 2024a) reveals layer-wise attention divergence - scattered patterns in lower layers gradually consolidate into instruction-critical focus in upper layers. This hierarchy informs its arithmetic-progressive cache allocation, preserving full instruction tokens while progressively reducing intermediate-layer retention.These methods demonstrate multidimensional approaches to token importance quantification. However, these eviction strategies inevitably induce attention distribution drift and error propagation, as permanently discarding low-salience tokens disrupts the intrinsic attention dynamics required for contextually coherent generation (Dao et al., 2022; Xiao et al., 2023b).
2.4 Cache merging
Token merging, initially pioneered in vision transformers (Bolya et al., 2022; Renggli et al., 2022), offers a viable pathway for KV cache optimization by fusing semantically similar tokens. The ToMe (Bolya et al., 2022) framework introduces training-free token fusion through similarity-based matching, enabling layer-wise token reduction without architectural modifications. In textual tasks, Dynamic Memory Compression (Zhang et al., 2024a) extends this concept by adaptively merging spatially proximate tokens, achieving substantial compression while retaining model efficacy through fine-tuning. Further innovations proposed in LESS (Dong et al., 2024) decouple computational bottlenecks by approximating attention mechanisms, where auxiliary networks compress low-priority tokens into compact representations with minimal impact on attention dynamics. From irreversible pruning to feature-preserving merging, from static heuristics to context-aware fusion policies, learnable mergers impose architectural complexity, demanding additional expertise in network tuning and increased computational resources, which complicates deployment.
3 Methodology
Building upon the conceptual foundation of ZeroMerge outlined earlier, this section delineates the technical framework underpinning its zero-shot compression mechanism. We propose a four-component methodology consisting of adaptive budget allocation, context-sensitive contribution evaluation, residual token merging, and stabilized attention projection. The interplay of these components ensures robust generalization across compression ratios and task domains without auxiliary parameters or fine-tuning, addressing limitations of prior eviction and merging strategies.
3.1 Preliminaries
Consider a transformer-based language model with $L$ layers, where each layer contains multi-head attention mechanisms. At decoding step $T$, let $\mathbf{K}_{T}\in\mathbb{R}^{T\times d}$ and $\mathbf{V}_{T}\in\mathbb{R}^{T\times d}$ denote the Key and Value cache matrices for a target attention head, formed by concatenating token-wise vectors up to the current step:
$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top},\quad\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top},$ (1)
where $\mathbf{k}_{t},\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the Key and Value vectors for the $t$-th token, and $d$ is the head dimension. For the query vector $\mathbf{q}_{T}\in\mathbb{R}^{d}$ at position $T$, the scaled dot-product attention computes output $\mathbf{o}_{i}$ via:
$a_{t}^{(T)}=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{s=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d})}.\quad\mathbf{o}^{(T)}=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}$ (2)
This formulation establishes the baseline for analyzing cache compression effects on attention distribution fidelity.
3.2 Budget Allocation
ZeroMerge strategically compresses the original $T$-length cache into a budget $B\ll T$ through tripartite allocation:
$B=B_{c}+B_{r}+B_{p},$ (3)
where $B_{c}$, $B_{r}$, and $B_{p}$ govern context preservation, residual absorption, and proximity maintenance, respectively. The compressed cache matrices aggregate three complementary components:
$\mathbf{K}_{B}=[\mathbf{K}_{c}\|\mathbf{K}_{r}\|\mathbf{K}_{p}],\quad\mathbf{V}_{B}=[\mathbf{V}_{c}\|\mathbf{V}_{r}\|\mathbf{V}_{p}],$ (4)
[…]
The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}\in\mathbb{R}^{T}$, which quantify contextual saliency. The proximity component $(\mathbf{K}_{p},\mathbf{V}_{p})$ preserves the latest $B_{p}$ tokens, capturing local attention patterns. The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through learned transformations. This configuration subsumes eviction-based methods when $B_{r}=0$, while enabling continuous interpolation between sparse and dense representations through budget reallocation
3.3 Contribution Evaluation
The context contribution metric $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$ dynamically quantifies each token’s cumulative influence across decoding steps. For the $t$-th token, its score $s_{t}^{(T)}$ evolves through Geometric decay integration of attention activations:
$s_{t}^{(T)}=\begin{cases}\lambda s_{t}^{(T-1)}+a_{t}^{(T)},&T>0\\ 0,&\text{otherwise}\end{cases},$ (5)
where $\lambda\in[0,1]$ is a momentum factor controlling historical preservation. This recursive formulation 1) attenuates early-stage attention signals through geometric decay, 2) maintains temporal locality by restricting updates to causally valid positions, and 3) enables online computation with $O(1)$ update cost per token.
3.4 Residual Token Merging
The residual component dynamically consolidates evicted tokens through similarity-driven aggregation. When merging a candidate token $(\mathbf{k}_{t},\mathbf{v}_{t})$ into $\mathbf{K}_{r}$, we:
  1. 1.
    Select the most compatible residual slot via maximum dot production:
    $\hat{r}=\mathop{\arg\max}\limits_{r\in\{1,\dots,B_{r}\}}\mathbf{k}_{r}^{\top}\mathbf{k}_{t}$ (6)
  2. 2.
    Update the selected slot through momentum-based fusion:
    $\mathbf{k}_{\hat{r}}\leftarrow\frac{w_{\hat{r}}\mathbf{k}_{\hat{r}}+\mathbf{k}_{t}}{w_{\hat{r}}+1},\quad\mathbf{v}_{\hat{r}}\leftarrow\frac{w_{\hat{r}}\mathbf{v}_{\hat{r}}+\mathbf{v}_{t}}{w_{\hat{r}}+1}$ (7)
  3. 3.
    Increment fusion count: $w_{\hat{r}}\leftarrow w_{\hat{r}}+1$
Figure 2 illustrates the dynamic cache evolution under budget parameters $B_{r}=2$, $B_{c}=4$, and $B_{p}=3$ during sequence length expansion from $T=12$ to $T=17$.
Figure 2: KV cache operation process with $B_{r}=2$, $B_{c}=4$, $B_{p}=3$, and sequence length expansion $T=12\rightarrow 17$: residual merging (2 slots), context preservation (top-4) and proximity maintenance (latest 3).
Algorithm 1 ZeroMerge Online Compression
1:  Input: Budgets $B_{c}$, $B_{r}$, $B_{p}$, decay $\lambda$
2:  Init: $\mathbf{K}_{B}\leftarrow(\emptyset,\emptyset,\emptyset)$, $\mathbf{V}_{B}\leftarrow(\emptyset,\emptyset,\emptyset)$
3:  for decoding step $T=1,2,\ldots$ do
4:     Compute attention scores $a_{T,t}$ via (2)
5:     Update contribution scores $\mathbf{s}^{(T)}$ via (5)
6:     $\mathbf{K}_{p}\leftarrow\mathbf{K}_{p}\cup\{\mathbf{k}_{T}\}$, $\mathbf{V}_{p}\leftarrow\mathbf{V}_{p}\cup\{\mathbf{v}_{T}\}$
7:     if $|\mathbf{K}_{p}|>B_{p}$ then
8:        Evict oldest token $(\mathbf{k}_{\text{old}},\mathbf{v}_{\text{old}})$ from $\mathbf{K}_{p}$, $\mathbf{V}_{p}$
9:        $\mathbf{K}_{c}\leftarrow\mathbf{K}_{c}\cup\{\mathbf{k}_{\text{old}}\}$, $\mathbf{V}_{c}\leftarrow\mathbf{V}_{c}\cup\{\mathbf{v}_{\text{old}}\}$
10:        if $|\mathbf{K}_{c}|>B_{c}$ then
11:           Evict lowest-score token $\hat{c}=\arg\min_{c}s_{c}^{(T)}$
12:           $\mathbf{K}_{r}\leftarrow\mathbf{K}_{r}\cup\{\mathbf{k}_{\hat{c}}\}$, $\mathbf{V}_{r}\leftarrow\mathbf{V}_{r}\cup\{\mathbf{v}_{\hat{c}}\}$
13:           if $|\mathbf{K}_{r}|>B_{r}$ then
14:              Merge via (6)–(7)
15:           end if
16:        end if
17:     end if
18:  end for
3.5 Attention Output Stabilization
Algorithm 1 progressively merges low-saliency tokens through dynamic cache updates. To faithfully reconstruct the original attention distribution from compressed representations, we introduce a compensated attention scoring mechanism. The revised attention computation evolves from Eq. 2 to:
$\hat{a}_{t}^{(T)}$ $=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{s=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d}+\alpha\log w_{s}\right)},$ (8)
$\mathbf{\hat{o}}^{(T)}$ $=\sum_{t=1}^{T}\hat{a}_{t}^{(T)}\mathbf{v}_{t},$
where $w_{j}$ denotes the fusion count of token $j$ (with $w_{j}=1$ for uncompressed tokens), and $\alpha\in(0,1]$ controls residual compensation intensity. This formulation addresses two key challenges:
  • Representation Bias Correction: Merged tokens aggregate multiple historical vectors through Eq. 7, creating a mismatch between their key vectors ($\mathbf{k}_{r}$) and the original value distribution. The $\log w_{j}$ term compensates for this representation shift.
  • Attention Mass Conservation: The compensation term preserves the relative attention mass between compressed and uncompressed tokens, preventing suppression of critical context.
Theorem 3.1.
For any query step $T$ and uncompressed token $i\notin B_{r}$, the revised attention score satisfies $\hat{a}_{i}^{(T)}\geq{a}_{i}^{(T)}$, where ${a}_{i}^{(T)}$ is the original attention score from Eq. 2.
Proof.
Let $r\in B_{r}$ be a residual slot merging $\{r_{1},...,r_{w_{r}}\}$ tokens. For compressed tokens, we establish an upper bound on their attention numerator:
$\text{num}(\hat{a}_{r}^{(T)})$ $=\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}+\alpha\log w_{r}\right)$ (9)
$\leq w_{r}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}\right)$
$=w_{r}\exp\left(\frac{1}{w_{r}}\sum_{m=1}^{w_{r}}\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right)$
$\leq\sum_{m=1}^{w_{r}}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right),$
where the first inequality uses $\alpha\leq 1$, and the second applies Jensen’s inequality to the convex exponential function. For uncompressed tokens ($w_{i}=1$), we derive:
$\hat{a}_{i}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t\notin B_{r}}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})+\sum_{r\in B_{r}}\text{num}(\hat{a}_{r}^{(T)})}$ (10)
$\geq\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}={a}_{i}^{(T)}.$
as the denominator contains compressed tokens’ upper-bounded contributions. ∎
This theoretical guarantee ensures uncompressed tokens maintain their relative attention dominance despite cache compression. The compensation mechanism effectively:
  • Preserves attention mass for critical context tokens
  • Prevents over-amplification of merged token scores
4 Experiment
In this chapter, we aim to verify the effectiveness of our proposed method in terms of performance preservation and memory reduction. Firstly, we introduce the experimental setup, including the selected model architectures, the baseline methods for comparison, and parameters such as the datasets used. Secondly, we verify the effects of our method on memory control and throughput. Moreover, we compare the generation quality of different methods in the field of long texts and tasks. Finally, we discuss the trade-off between memory, time, and performance as well as the impact of relevant hyperparameters on the model.
4.1 Experiment setup
The backbone models selected in the experiment are some small-scale open-sourced LLMs, including LLaMA-2 7B, LLaMa3-8B-Instruct, and Mistral-7B-Instruct. The experimental equipment is an NVIDIA A800-80GB GPU. For the dataset, we choose LongBench, which is a comprehensive benchmark dataset specifically designed for evaluating the performance of different methods on tasks that involve long-context inputs. It contains a wide variety of natural language processing tasks with different lengths of text sequences, covering multiple domains such as news, novels, and academic papers, thus providing a rich and diverse set of test cases for our experiments.
The baselines for comparison are different KV cache management methods. Specifically, they include:
  • Full KV (FKV): Caches all keys and values per token in each layer, serving as the vanilla baseline with maximal computational load but theoretically optimal performance.
  • Heavy Hitter Oracle (H2O): Dynamically manages KV cache using cumulative attention thresholds to balance recent tokens and Heavy Hitters (H2). Leverages layer-wise attention score averaging across queries to identify critical tokens, achieving adaptive compression while maintaining task-specific semantics.
  • LESS: Introduces recurrent KV state synthesis through layer-aware dynamic merging. Combines temporal locality exploitation (retaining recent tokens) with spatial compression via parameterized merging operators. Implements attention rectification to compensate for distribution shifts caused by merged states, enabling sublinear cache growth while preserving contextual coherence.
  • ZeroMerge (Ours): Builds upon merging paradigm but eliminates parametric dependencies through heuristic-based 0-shot merging. Strictly enforces preset cache bounds via two-phase compression: 1) Attention sink preservation for initial/recent tokens, 2) Distance-aware similarity clustering for intermediate tokens. Minimizes embedding distortion through revised attention scoring that accounts for merged token positional densities, achieving better generalization across compression ratios than prior learning-based methods.
4.2 Inference efficiency and resource consumption
Figure 3: VRAM Utilization and Decoding Throughput Across Sequence Lengths. ZeroMerge enforces constant memory footprint (43GB) and sustains 9 tokens/sec decoding rate beyond 54K tokens, eliminating OOM via dynamic KV cache compression.
We evaluate the long - context processing efficiency on a single NVIDIA A800 GPU with 80GB VRAM (Video Random Access Memory). The experiment uses the LLaMA2 - 7B model, which has 32 layers, 32 attention heads, and an embedding dimension of 4096, in float32 precision. The benchmark is composed of a synthetic QA task, where the cache tokens range from 20K to 54K. For the baseline measurements, we use full KV caching without compression. In contrast, the ZeroMerge method enforces a token cache budget of 18K.
The original Llama2-7B model (float32) required 25GB VRAM for parameter loading alone. During prefill phase, processing 20K prompt tokens generated 20GB KV cache across 32 layers, totally consuming 56% of total available VRAM (Fig.3). In baseline inference, KV cache grew linearly with sequence length at 1MB/token. By 54K tokens, cache memory dominated the mast of VRAM, triggering OOM errors. Concurrently, decoding throughput degraded from 9 tokens/sec (20K context) to 4 tokens/sec (52K context)—due to the increased computation of attention score.
By enforcing a 18K token cache ceiling through two-phase compression (attention sink preservation + similarity-based merging), ZeroMerge: Reduced prefill-phase KV cache to 18GB (10% reduction); Maintained constant 43GB VRAM usage during decoding (vs. baseline’s 45 to 79 GB linear growth); Sustained stable throughput at 9 tokens/sec across 20K–54K token contexts; Eliminated OOM constraints through dynamic cache compression.
4.3 Workload-Scalable KV Efficiency
We evaluate ZeroMerge against three baseline methods (FullKV, H2O, LESS) on an NVIDIA A800 GPU (80GB VRAM) using LLaMA2-7B (float32) with diverse workloads: sequence lengths (2K–54K tokens), batch sizes (4–16), and document-QA tasks. Each method is tested under identical decoding constraints (temperature=0.0, greedy sampling) to isolate computational overhead. Metrics include throughput (tokens/sec), latency (sec), and OOM thresholds.
Table 1: Workload-Scalable KV Cache Compression: ZeroMerge Outperforms Baselines in Throughput (tokens/sec) and Latency (seconds) Across Sequence Lengths and Batch Sizes.
Seq.length Batch size Model size throughput(tokens/S) $\Uparrow$ / latency(S) $\Downarrow$
FullKV H2O (5%) Less (5%) ZeroMerge (5%)
4096+4096 4 7B 62.4 / 262.1 43.9 / 372.5 15.0 / 1086.2 77.0 / 212.7
4096+4096 8 7B 65.5 / 500.3 OOM OOM 146.7 / 223.2
2048+2048 16 7B 133.1 / 246.2 86.1 / 380.1 OOM 281.9 / 178.36
4096+4096 16 13B OOM - - 178.2 / 367.6
As demonstrated in Table 1, ZeroMerge consistently outperforms existing methods across diverse workload configurations. At moderate batch sizes (batch=4, 4K+4K tokens), ZeroMerge achieves 23.4% higher throughput (77.0 vs 62.4 tokens/sec) and 18.9% lower latency (212.7s vs 262.1s) compared to FullKV, while H2O and LESS suffer 29.7% and 75.9% performance degradation respectively. Under high-throughput demands (batch=8, 4K+4K tokens), ZeroMerge successfully processes requests where both H2O and LESS encounter OOM errors, delivering 124% throughput improvement (146.7 vs 65.5 tokens/sec) over FullKV. Most notably, at large batch sizes (batch=16, 2K+2K tokens), ZeroMerge sustains 2.1× higher throughput (281.9 vs 133.1 tokens/sec) with 52.8% latency reduction (116.2s vs 246.2s), showcasing its unique capability to decouple computational load from sequence length. These results validate ZeroMerge’s dual advantage, eliminating memory bottlenecks through strict cache budgeting.
4.4 Inference quality
We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics. The study compares three KV cache compression methods across two 7B models (Llama2, Falcon) at 20%, 10%, and 5% cache budgets. H2O balances recent tokens and heavy hitters per its original design, while LESS employs its recommended C4-trained merging network. ZeroMerge operates with $\alpha=0.6$ (influence ratio of residual cache ) and $\lambda=0.98$ (the decay rate of the historical contribution) without task-specific training.
Our evaluation demonstrates that ZeroMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods. When compressing the KV cache to 5% of its original size, ZeroMerge maintains near-original performance on Llama2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions. This advantage becomes more pronounced at higher compression ratios, where ZeroMerge’s performance preservation significantly outperforms both baselines.
On Falcon-7B, a model with distinct architectural characteristics, ZeroMerge demonstrates stronger generalization than parameter-dependent approaches. While LESS suffers severe quality degradation—likely due to its reliance on C4 dataset training—ZeroMerge retains over half of the baseline performance even at extreme 5% compression. In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
The results highlight two critical trends: First, higher compression ratios amplify ZeroMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies. Second, architecture-agnostic design enables consistent cross-model robustness, avoiding the domain adaptation challenges faced by data-driven methods like LESS.
Table 2: Model-Agnostic Text Generation Quality under KV Cache Compression: ROUGE-1/2/L Scores for Llama2-7B and Falcon-7B Across 20%, 10%, and 5% Cache Budgets
method Llama2 7B Falcon 7B
FULL 30.59 / 11.34 / 25.50 27.06 / 8.79 / 22.39
20% H2O 30.83 / 11.43 / 25.71 24.18 / 7.47 / 20.14
20% LESS 30.47 / 11.23 / 25.44 24.90 / 7.97 / 20.76
20% ZeroMerge 31.62 / 12.40 / 26.42 25.19 / 8.69 / 21.34
10% H2O 30.18 / 11.32 / 25.28 13.03 / 3.16 / 11.02
10% LESS 30.74 / 11.22 / 25.58 8.99 / 2.55 / 7.50
10% ZeroMerge 31.83 / 12.47 / 26.75 20.92 / 6.74 / 17.50
5% H2O 28.92 / 10.81 / 24.35 12.02 / 2.08 / 10.36
5% LESS 29.98 / 11.09 / 25.02 7.75 / 1.15 / 6.76
5% ZeroMerge 30.60 / 11.67 / 25.72 15.04 / 3.29 / 12.73
Acknowledgements
Do not include acknowledgements in the initial version of the paper submitted for blind review.
If a paper is accepted, the final camera-ready version can (and usually should) include acknowledgements. Such acknowledgements should be placed at the end of the section, in an unnumbered section that does not count towards the paper page limit. Typically, this will include thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
Impact Statement
Authors are required to include a statement of the potential broader impact of their work, including its ethical aspects and future societal consequences. This statement should be in an unnumbered section at the end of the paper (co-located with Acknowledgements – the two may appear in either order, but both must be before References), and does not count toward the paper page limit. In many cases, where the ethical impacts and expected societal implications are those that are well established when advancing the field of Machine Learning, substantial discussion is not required, and a simple statement such as the following will suffice:
“This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.”
The above statement can be used verbatim in such cases, but we encourage authors to think about whether there is content which does warrant further discussion, as this statement will be apparent if the paper is later flagged for ethics review.

### v2

@@
-ZeroMerge: Parameter-Free KV Cache Compression for Memory-Efficient Long-Context LLMs
+ZSMerge: Zero-Shot KV Cache Compression for Memory-Efficient Long-Context LLMs
 Abstract
-The linear growth of key-value (KV) cache memory and quadratic computational complexity pose significant bottlenecks for large language models (LLMs) in long-context processing.
-While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often suffer from irreversible information loss or require costly parameter retraining.
-We propose ZeroMerge, a dynamic zero-shot compression framework that achieves efficient cache management through three key innovations: (1) Fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) A residual merging mechanism that preserves critical context through compensated attention scoring, and (3) Parameter-free adaptation compatible with diverse LLM architectures without retraining.
-Comprehensive evaluations across LLaMA-2 model demonstrate that ZeroMerge maintains full-cache performance at 5% compression ratios while doubling inference throughput at 40K token lengths.
-The method effectively balances memory efficiency, generation quality, and deployment flexibility, advancing practical long-context LLM applications.
-The code is available at https://github.com/SusCom-Lab/ZeroMerge.
+The linear growth of key-value (KV) cache memory and quadratic computational complexity in attention mechanisms pose significant bottlenecks for large language models (LLMs) in long-context processing.
+While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often incur irreversible information loss or require costly parameter retraining.
+To this end, we propose ZSMerge, a dynamic KV cache compression framework designed for efficient cache management, featuring three key operations: (1) fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) a residual merging mechanism that preserves critical context through compensated attention scoring, and (3) a zero-shot adaptation mechanism compatible with diverse LLM architectures without requiring retraining.
+ZSMerge significantly enhances memory efficiency and inference speed with negligible performance degradation across LLMs.
+When applied to LLaMA2-7B, it demonstrates a 20:1 compression ratio for key-value cache retention (reducing memory footprint to 5% of baseline) while sustaining comparable generation quality, coupled with triple throughput gains at extreme 54k-token contexts that eliminate out-of-memory failures.
+The code is available at https://github.com/SusCom-Lab/ZSMerge.
 Machine Learning, ICML
 1 Introduction
-The advancement of the Transformer architecture has brought about revolutionary changes in handling sequence data, with large language models (LLMs) being a prime illustration.
-LLMs have achieved significant progress, surpassing humans in diverse applications.
-However, handling long contexts has posed challenges, particularly the linear expansion of the KV cache in proportion to the sequence length.
-For instance, the deployment of a Llama 2 7B model consumes 26GB of memory.
-When the token length reaches 32K, the 32G of memory occupied by the KV cache becomes the main memory consumer.
-This, along with the quadratic computational complexity, has impeded LLM integration and performance, spurring research on optimizing the KV cache for enhanced inference efficiency.
-Abundant innovative studies on architecture compression alleviate the shortage of memory (Sarah et al., 2024; Ashkboos et al., 2024).
-Parameter quantization techniques(Dettmers et al., 2022; Xiao et al., 2023a) further reduce memory footprint by representing KV cache with low-bit precision (e.g., 4-bit integers), though at the cost of numerical accuracy preservation.
-More efforts focus on the management of the KV cache to directly address this challenge.
-Some sparse methods(Zhang et al., 2024b, a) reserve partial top-important tokens and eliminate others to mitigate the harm to the original generation quality.
-These selection criteria are based on the historical contribution in the decoding procedure, spatio-temporal relationship, or data distribution.
-One-sided and short-sighted pre-pruning strategies inevitably disrupt the distribution of KV embeddings and introduce unrecoverable information loss.
-Hybrid approaches combining quantization and sparsity (Dettmers et al., 2023) attempt to balance accuracy and compression ratio, yet still require delicate bit allocation strategies.
-Low rank compression techniques are utilized to correct the systematic drawback in some works (Dong et al., 2024).
-However, the newly introduced architectures or the retraining requirement of the model parameters constitutes a significant burden.
-One desirable KV management strategy should 1) strictly control the memory consumption to break the context length limits 2) minimize the generation quality degradation 3) be easily applied to various LLM backbones without extra refinement cost.
-Motivated by these goals, we proposed ZeroMerge, a more efficient and comprehensive dynamic zero-shot KV management algorithm.
-  1. 1.
+The advancement of the Transformer architecture has revolutionized sequence data processing, with Large Language Models (LLMs) being a prime illustration.
+LLMs have achieved remarkable progress, surpassing human capabilities in diverse applications.
+While multi-turn dialogue and long-context understanding are critical scenarios, handling extended sequences remains challenging due to the linear expansion of the Key-Value (KV) cache (which stores intermediate attention keys and values during generation to avoid redundant computations) in proportion to text length.
+For instance, deploying a LLaMA2-7B model will consume about 26GB of GPU memory; when the token length reaches 32K, the KV cache alone occupies 32G of memory, becoming the primary memory consumption.
+This, along with the quadratic computational complexity, significantly restricts LLM integration and performance, thus motivating recent research efforts toward optimizing KV cache utilization for enhanced inference efficiency.
+To mitigate the large memory footprint of the KV cache, various optimization strategies have emerged.
+Some modify the core model architecture (Sarah et al., 2024; Ashkboos et al., 2024) or employ quantization techniques for lower precision representation (Dettmers et al., 2022; Xiao et al., 2023a).
+Other methods directly target the KV cache bottleneck during inference using context-aware techniques.
+Common among these is sparse approximation or token eviction, retaining only crucial tokens while discarding others (Zhang et al., 2024b, a); however, this can disrupt KV embeddings and cause information loss.
+To mitigate this, alternative context-aware techniques merge or fuse information from similar tokens (Dong et al., 2024), often using low-rank compression, though these may introduce architectural overhead or require model retraining.
+Ideally, an effective and practical KV cache management strategy should 1) strictly control the memory consumption to break the context length limits, 2) minimize the generation quality degradation, and 3) readily applicable across diverse LLM backbones without necessitating fine-tuning or architectural changes (zero-shot compatibility).
+Motivated by these goals, we propose ZSMerge, an efficient, comprehensive, and dynamic zero-shot KV management algorithm, with the following key features:
   • Fine-grained memory allocation.
-ZeroMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and its own data distribution characteristics to conduct fine-grained memory management at the head level, making the method more adaptable to different types of long text tasks.
-It has achieved superior generation quality compared to the sparse strategy and is comparable to the retrained model on models.
-The performance of 5% of the KV cache bugdet is still on par with the complete model.
-  2. 2.
-    Low-cost adaptation and deployment.
-ZeroMerge does not introduce any model parameters and does not require retraining or model fine-tuning to adapt to different compression ratios or various task types.
-It is adapted to various mainstream model architectures such as Llama, Falcon, and Mistral.
-  3. 3.
-    Efficient compression strategy.
-ZeroMerge adopts a light weight operational overhead of linear operational complexity and achieves more than twice the throughput of the original model at a token length of 50K.
-We have proved that the cache compression mechanism adopted by ZeroMerge does not drown out the original contribution of those intact tokens as the processing sequence grows.
+ZSMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and its own data distribution characteristics to conduct fine-grained, head-level memory management.
+This adaptability allows it to maintain superior generation quality compared to the sparse strategies and achieve performance comparable to retraining-dependent approaches, even when operating with significantly reduced memory budgets.
+This ensures effectiveness across diverse long-context tasks.
+  • Zero-shot and low-cost Integration.
+Designed for ease of use, ZSMerge does not introduce any model parameters and thus requires no retraining or model fine-tuning to adapt to different compression ratios or various task types.
+This also makes it compatible to various mainstream LLM architectures such as LLaMA, Falcon, and Mistral, thereby minimizing deployment overhead.
+  • Efficient and throughput-boosting strategy.
+Characterized by linear operational complexity, ZSMerge’s compression mechanism is computationally lightweight.
+This efficiency brings significant throughput improvement, e.g., it helps achieve over triple the inference throughput compared to the original model at a token length of 54K.
+Furthermore, we have proved that the KV cache compressed by ZSMerge can effectively preserves the informational contribution of retained tokens, preventing signal degradation even as context length increases substantially.
 2 Related Work
-Figure 1: Comparative analysis of KV cache optimization strategies: (a) Model architecture compression reduces both model parameters and KV cache size via multi-query attention heads; (b) Parameter/data quantization trades numerical precision for cache footprint reduction through low-bit representations; (c) Sparse approximation directly addresses memory-compute growth by evicting low-contribution tokens, albeit inducing attention distribution drift; (d) Token merge mitigates sparse caching errors by learning residuals between full and sparse attention outputs, yet introduces overhead from auxiliary networks or fine-tuning; (e) ZeroMerge (proposed) achieves zero-shot compression competitive with sparse+residual methods without added parameters, while demonstrating superior generalization across task domains and compression ratios.
-The phenomenon of emergence in large-scale models has ignited new enthusiasm and provided fresh directions for exploring the nonlinear leaps in reasoning and generative capabilities of LLMs—driven by the expansion of model scale and complexity.
-The unprecedented parameter sizes and intensive computational demands have endowed LLMs with remarkable reasoning and generative abilities, yet they also pose significant challenges, hindering their cost-effective and large-scale deployment across various scenarios.
-Numerous efforts have been made to address the issue.
-2.1 Model architecture compression
-Works such as MiniLLM (Gu et al., 2024) and SliceGPT (Ashkboos et al., 2024) have been specifically devoted to ensure that the performance of sample tasks remains intact for more compact models.
-MiniLLM focuses on knowledge distillation of large language models, aiming to transfer the knowledge from large and powerful models to smaller ones, thereby reducing the number of parameters in the latter.
-On the other hand, SliceGPT adopts a different approach by replacing weight matrix with a smaller dense matrix.
-Through these innovative methods, they both strive to strike a balance between controlling the parameter count and maintaining satisfactory performance on relevant tasks.
-In the research of attention mechanisms, (Voita et al., 2019) pointed out that the contributions of different attention heads to the overall performance of the model in the multi-head self-attention mechanism are not equal.
-It also provided a novel pruning method at the attention head level.
-Similarly, a large number of works related to Multi-Query Attention (MQA) enable different attention “head” to share a set of “keys” and “values”, thus significantly reducing the size of these tensors and the memory bandwidth requirements for incremental decoding (Shazeer, 2019; Ainslie et al., 2023).
-2.2 Parameter and data quantization
-Quantization techniques have emerged as a crucial approach to address the computational and storage challenges of large language models.
-GPTQ(Frantar et al., 2022) is a one-shot weight quantization method based on approximate second-order information.
-It can quantize large GPT models with a significant number of parameters in a relatively short time and with minimal accuracy degradation.
-This enables models to be executed on a single GPU for generative inference and provides end-to-end inference speedups when using different GPUs.
-INT-FlashAttention(Chen et al., 2024) is the first INT8 quantization architecture compatible with the forward workflow of FlashAttention.
-It significantly improves the inference speed of FlashAttention on Ampere GPUs by implementing fully INT8 activations and general matrix-multiplication kernels.
-It also serves as a general token-level post-training quantization framework compatible with other data formats.
-They not only reduce the model size but also enhance the inference speed and manage quantization errors, making large language models more accessible and efficient for various applications.
-2.3 Sparse approximation
-Other studies directly focus on the KV cache, another performance bottleneck induced by long sequence tasks, striving to fundamentally address the issues of storage growth and quadratic computational complexity that accompany the increase in sequence sample length.
-The universality of sparse attention values (Ji et al., 2021; Deng et al., 2024) has provided a new direction for improving the computational efficiency and reducing the memory footprint of models.
-(Roy et al., 2021) dynamically routes the input tokens to different attention heads based on their content, enabling each attention head to focus only on the relevant parts, thus enhancing the computational efficiency while maintaining the model’s performance.
-(Wang et al., 2020) takes advantage of the low rank property and projects it into a low-order matrix, thereby avoiding the attention layer with quadratic computational complexity.
-Another token aware sparse approach, also known as eviction strategy, employ diverse token importance metrics to address memory bottlenecks.
-The foundational work by (Xiao et al., 2023b) identifies the attention sink phenomenon, where initial tokens disproportionately attract attention scores regardless of positional distance.
-Building on this insight, StreamingLLM combines a sliding temporal window with persistent retention of initial token KV states, enabling stable infinite-length generation.
-Similarly, LM-infinite (Han et al., 2024) introduces Lambda-shaped attention masks to differentially strengthen local (recent tokens) and global (starting tokens) semantic modeling.
-H2O (Zhang et al., 2024b)proposes Heavy Hitter token detection through cumulative attention score thresholds, dynamically evicting non-critical tokens during decoding.
-Simultaneously, PyramidKV(Zhang et al., 2024a) reveals layer-wise attention divergence - scattered patterns in lower layers gradually consolidate into instruction-critical focus in upper layers.
-This hierarchy informs its arithmetic-progressive cache allocation, preserving full instruction tokens while progressively reducing intermediate-layer retention.These methods demonstrate multidimensional approaches to token importance quantification.
-However, these eviction strategies inevitably induce attention distribution drift and error propagation, as permanently discarding low-salience tokens disrupts the intrinsic attention dynamics required for contextually coherent generation (Dao et al., 2022; Xiao et al., 2023b).
-2.4 Cache merging
-Token merging, initially pioneered in vision transformers (Bolya et al., 2022; Renggli et al., 2022), offers a viable pathway for KV cache optimization by fusing semantically similar tokens.
-The ToMe (Bolya et al., 2022) framework introduces training-free token fusion through similarity-based matching, enabling layer-wise token reduction without architectural modifications.
-In textual tasks, Dynamic Memory Compression (Zhang et al., 2024a) extends this concept by adaptively merging spatially proximate tokens, achieving substantial compression while retaining model efficacy through fine-tuning.
-Further innovations proposed in LESS (Dong et al., 2024) decouple computational bottlenecks by approximating attention mechanisms, where auxiliary networks compress low-priority tokens into compact representations with minimal impact on attention dynamics.
-From irreversible pruning to feature-preserving merging, from static heuristics to context-aware fusion policies, learnable mergers impose architectural complexity, demanding additional expertise in network tuning and increased computational resources, which complicates deployment.
-3 Methodology
-Building upon the conceptual foundation of ZeroMerge outlined earlier, this section delineates the technical framework underpinning its zero-shot compression mechanism.
+Figure 1: Comparative analysis of KV cache optimization strategies: (A) Sparse approximation directly addresses memory-compute growth by evicting low-contribution tokens, albeit inducing attention distribution drift.
+(B) Token merge mitigates sparse caching errors, often by learning residuals implicitly or explicitly, yet introduces overhead from auxiliary networks or fine-tuning.
+(C) ZSMerge (proposed) achieves zero-shot compression competitive with methods requiring learned components, doing so without added parameters while demonstrating superior generalization.
+The escalating memory and computational demands of large language models (LLMs), particularly from KV cache growth in long-context scenarios, have spurred diverse optimization strategies.
+These are broadly categorized into context-agnostic and contextually adaptive approaches.
+2.1 Context-Agnostic Optimization
+Context-agnostic methods universally reduce resource consumption by altering the model’s structure or numerical representation, irrespective of the input sequence.
+Foundational techniques include Structural Compression via knowledge distillation (e.g., MiniLLM (Gu et al., 2024)) or matrix factorization (e.g., SliceGPT (Ashkboos et al., 2024)).
+Attention Mechanism Optimization, such as head pruning (Voita et al., 2019) and key-value sharing in Multi-Query (MQA (Shazeer, 2019)) and Grouped-Query Attention (GQA (Ainslie et al., 2023)), significantly cuts memory bandwidth.
+Simultaneously, Numerical Optimization tackles precision redundancy through techniques like post-training quantization (e.g., GPTQ (Frantar et al., 2022)), hardware-aligned quantization (e.g., INT-FlashAttention (Chen et al., 2024)), and exploiting low-rank properties (e.g., Linformer (Wang et al., 2020)).
+These strategies collectively alleviate memory bottlenecks via architectural simplification and precision reduction.
+2.2 Contextually Adaptive Optimization
+While context-agnostic methods offer baseline efficiency, recent research increasingly aims to directly target the KV cache—a critical bottleneck in long-sequence tasks driven by storage growth and quadratic computational complexity.
+The observed sparsity in attention values (Ji et al., 2021; Deng et al., 2024) motivates dynamic, context-aware strategies.
+Figure 1 illustrates the main categories of these adaptive approaches:
+(A) Sparse Approximation (Eviction): This class directly addresses memory-compute growth by evicting low-contribution tokens, based on the premise that not all history is equally vital.
+Work like (Xiao et al., 2023b) identified ”attention sinks,” leading to methods like StreamingLLM (Xiao et al., 2023b) which retains initial tokens alongside a sliding window.
+LM-infinite (Han et al., 2024) uses specialized masks to balance local and global context.
+H2O (Zhang et al., 2024b) employs attention score thresholds for dynamic eviction, while PyramidKV (Zhang et al., 2024a) uses layer-wise attention patterns for progressive caching.
+Dynamic routing (Roy et al., 2021) also fits here.
+However, eviction strategies inevitably induce attention distribution drift and error propagation by permanently discarding tokens (Dao et al., 2022; Xiao et al., 2023b).
+(B) Token Merging: To mitigate information loss from eviction, token merging fuses similar tokens instead of discarding them.
+Pioneered in vision (e.g., ToMe (Bolya et al., 2022)), it applies to text as well.
+Dynamic Memory Compression adaptively merges spatially proximate tokens, often requiring fine-tuning.
+LESS (Dong et al., 2024) uses auxiliary networks to compress tokens with minimal attention impact.
+While merging mitigates sparse caching errors by learning residuals (implicitly or explicitly), it often introduces overhead from auxiliary networks or fine-tuning, complicating deployment.
+(C) ZSMerge (Proposed): Addressing the limitations of both prior approaches, we propose ZSMerge.
+It achieves zero-shot compression competitive with sparse+residual methods without added parameters, while demonstrating superior generalization across task domains and compression ratios.
+ZSMerge offers a parameter-free, tuning-free approach, balancing efficiency and generation performance.
+3 ZSMerge Methodology
+Building upon the conceptual foundation of ZSMerge outlined earlier, this section delineates the technical framework underpinning its zero-shot compression mechanism.
 We propose a four-component methodology consisting of adaptive budget allocation, context-sensitive contribution evaluation, residual token merging, and stabilized attention projection.
@@
 3.1 Preliminaries
-Consider a transformer-based language model with $L$ layers, where each layer contains multi-head attention mechanisms.
-At decoding step $T$, let $\mathbf{K}_{T}\in\mathbb{R}^{T\times d}$ and $\mathbf{V}_{T}\in\mathbb{R}^{T\times d}$ denote the Key and Value cache matrices for a target attention head, formed by concatenating token-wise vectors up to the current step:
-$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top},\quad\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top},$ (1)
-where $\mathbf{k}_{t},\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the Key and Value vectors for the $t$-th token, and $d$ is the head dimension.
-For the query vector $\mathbf{q}_{T}\in\mathbb{R}^{d}$ at position $T$, the scaled dot-product attention computes output $\mathbf{o}_{i}$ via:
-$a_{t}^{(T)}=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{s=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d})}.\quad\mathbf{o}^{(T)}=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}$ (2)
+Consider an $L$-layer transformer with multi-head attention mechanisms. For a target attention head at decoding step $T$, the cached Key and Value matrices are defined as:
+$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top}\in\mathbb{R}^{T\times d},$ (1)
+$\quad\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top}\in\mathbb{R}^{T\times d},$
+where $\mathbf{k}_{t},\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the Key/Value vectors for the $t$-th token.
+For the query vector $\mathbf{q}_{T}\in\mathbb{R}^{d}$ at position $T$, the scaled dot-product attention computes output $\mathbf{o}^{(T)}$ via:
+$a_{t}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{s=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d})},$ (2)
+$\quad\mathbf{o}^{(T)}$ $=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}$
 This formulation establishes the baseline for analyzing cache compression effects on attention distribution fidelity.
 3.2 Budget Allocation
-ZeroMerge strategically compresses the original $T$-length cache into a budget $B\ll T$ through tripartite allocation:
-$B=B_{c}+B_{r}+B_{p},$ (3)
-where $B_{c}$, $B_{r}$, and $B_{p}$ govern context preservation, residual absorption, and proximity maintenance, respectively. The compressed cache matrices aggregate three complementary components:
-$\mathbf{K}_{B}=[\mathbf{K}_{c}\|\mathbf{K}_{r}\|\mathbf{K}_{p}],\quad\mathbf{V}_{B}=[\mathbf{V}_{c}\|\mathbf{V}_{r}\|\mathbf{V}_{p}],$ (4)
+ZSMerge strategically compresses the original $T$-length cache into a budget $B\ll T$ through tripartite allocation:
+$B=B_{p}+B_{c}+B_{r},$ (3)
+where $B_{p}$, $B_{c}$, and $B_{r}$ govern context preservation, residual absorption, and proximity maintenance, respectively. The compressed cache matrices aggregate three complementary components:
+$\mathbf{K}_{B}=[\mathbf{K}_{p}\|\mathbf{K}_{c}\|\mathbf{K}_{r}],\quad\mathbf{V}_{B}=[\mathbf{V}_{p}\|\mathbf{V}_{c}\|\mathbf{V}_{r}],$ (4)
 […]
-The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}\in\mathbb{R}^{T}$, which quantify contextual saliency.
 The proximity component $(\mathbf{K}_{p},\mathbf{V}_{p})$ preserves the latest $B_{p}$ tokens, capturing local attention patterns.
+The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$, which quantify contextual saliency.
 The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through learned transformations.
-This configuration subsumes eviction-based methods when $B_{r}=0$, while enabling continuous interpolation between sparse and dense representations through budget reallocation
+This configuration subsumes eviction-based methods when $B_{r}=0$, while enabling continuous interpolation between sparse and dense representations through budget reallocation.
 3.3 Contribution Evaluation
-The context contribution metric $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$ dynamically quantifies each token’s cumulative influence across decoding steps.
-For the $t$-th token, its score $s_{t}^{(T)}$ evolves through Geometric decay integration of attention activations:
+The contribution scores $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$ dynamically quantify each token’s cumulative influence across decoding steps.
+For the $t$-th token, its score $s_{t}^{(T)}$ evolves through exponential decay integration of attention activations:
 $s_{t}^{(T)}=\begin{cases}\lambda s_{t}^{(T-1)}+a_{t}^{(T)},&T>0\\ 0,&\text{otherwise}\end{cases},$ (5)
-where $\lambda\in[0,1]$ is a momentum factor controlling historical preservation.
-This recursive formulation 1) attenuates early-stage attention signals through geometric decay, 2) maintains temporal locality by restricting updates to causally valid positions, and 3) enables online computation with $O(1)$ update cost per token.
+where the decay factor $\lambda\in[0,1]$ operates as a temporal discounting parameter analogous to reinforcement learning credit assignment, governing the exponential decay of historical attention contributions.
+We recommend setting it to a relatively large value to balance the contributions of long-term and short-term attention scores.
+Throughout this paper, $\lambda$ is configured as $0.98$ unless otherwise specified, as scores from 300 steps earlier become attenuated to approximately 0.002 through exponential decay ($0.98^{300}\approx 0.002$).
 3.4 Residual Token Merging
-The residual component dynamically consolidates evicted tokens through similarity-driven aggregation. When merging a candidate token $(\mathbf{k}_{t},\mathbf{v}_{t})$ into $\mathbf{K}_{r}$, we:
+The residual component dynamically consolidates evicted tokens through similarity-driven aggregation.
+When merging a candidate token $(\mathbf{k}_{t},\mathbf{v}_{t})$ into $\mathbf{K}_{r}$, we adopt the following three-step procedure:
   1. 1.
@@
 Figure 2: KV cache operation process with $B_{r}=2$, $B_{c}=4$, $B_{p}=3$, and sequence length expansion $T=12\rightarrow 17$: residual merging (2 slots), context preservation (top-4) and proximity maintenance (latest 3).
-Algorithm 1 ZeroMerge Online Compression
-1:  Input: Budgets $B_{c}$, $B_{r}$, $B_{p}$, decay $\lambda$
+Algorithm 1 ZSMerge Online Compression
+1:  Input: Budgets $B_{p}$, $B_{c}$, $B_{r}$, decay $\lambda$
 2:  Init: $\mathbf{K}_{B}\leftarrow(\emptyset,\emptyset,\emptyset)$, $\mathbf{V}_{B}\leftarrow(\emptyset,\emptyset,\emptyset)$
@@
 11:           Evict lowest-score token $\hat{c}=\arg\min_{c}s_{c}^{(T)}$
-12:           $\mathbf{K}_{r}\leftarrow\mathbf{K}_{r}\cup\{\mathbf{k}_{\hat{c}}\}$, $\mathbf{V}_{r}\leftarrow\mathbf{V}_{r}\cup\{\mathbf{v}_{\hat{c}}\}$
-13:           if $|\mathbf{K}_{r}|>B_{r}$ then
-14:              Merge via (6)–(7)
-15:           end if
+12:           if $|\mathbf{K}_{r}|+1\leq B_{r}$ then
+13:              $\mathbf{K}_{r}\leftarrow\mathbf{K}_{r}\cup\{\mathbf{k}_{\hat{c}}\}$, $\mathbf{V}_{r}\leftarrow\mathbf{V}_{r}\cup\{\mathbf{v}_{\hat{c}}\}$
+14:           else
+15:              Merge token $(\mathbf{k}_{\hat{c}},\mathbf{v}_{\hat{c}})$ into $\mathbf{K}_{r}$ via (6)-(7)
 16:           end if
 17:        end if
-18:  end for
+18:     end if
+19:  end for
 3.5 Attention Output Stabilization
@@
 2 to:
-$\hat{a}_{t}^{(T)}$ $=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{s=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d}+\alpha\log w_{s}\right)},$ (8)
+$\hat{a}_{t}^{(T)}$ $=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{i=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d}+\alpha\log w_{i}\right)},$ (8)
 $\mathbf{\hat{o}}^{(T)}$ $=\sum_{t=1}^{T}\hat{a}_{t}^{(T)}\mathbf{v}_{t},$
-where $w_{j}$ denotes the fusion count of token $j$ (with $w_{j}=1$ for uncompressed tokens), and $\alpha\in(0,1]$ controls residual compensation intensity.
-This formulation addresses two key challenges:
-  • Representation Bias Correction: Merged tokens aggregate multiple historical vectors through Eq.
+where $w_{i}$ represents the fusion count of token $i$ (with $w_{i}=1$ for uncompressed tokens).
+The hyperparameter $\alpha\in(0,1]$ regulates residual compensation intensity, with empirical analysis suggesting an optimal range of $\alpha\in[0.6,0.8]$ for balanced performance.
+Notably, setting $\alpha=0$ reduces the merging algorithm to a degenerate case equivalent to a pure eviction policy, eliminating residual compensation entirely.
+With Eq. 8, we address two key challenges:
+  • Representation Bias Correction: Merged tokens aggregate multiple historical vectors via Eq.
 7, creating a mismatch between their key vectors ($\mathbf{k}_{r}$) and the original value distribution.
 The $\log w_{j}$ term compensates for this representation shift.
-  • Attention Mass Conservation: The compensation term preserves the relative attention mass between compressed and uncompressed tokens, preventing suppression of critical context.
+  • Attention Mass Conservation: The compensation term preserves the relative attention mass between compressed and uncompressed tokens, ensuring residual compensation does not suppress critical uncompressed components.
 Theorem 3.1.
@@
 as the denominator contains compressed tokens’ upper-bounded contributions. ∎
-This theoretical guarantee ensures uncompressed tokens maintain their relative attention dominance despite cache compression. The compensation mechanism effectively:
-  • Preserves attention mass for critical context tokens
-  • Prevents over-amplification of merged token scores
-4 Experiment
+This theoretical guarantee ensures that uncompressed tokens retain their relative dominance in attention allocation despite cache compression, even as the denominator accounts for upper-bounded contributions from compressed tokens.
+The compensation mechanism effectively preserves attention mass for critical context tokens while preventing over-amplification of merged token scores, thereby maintaining the integrity of the original attention distribution under compression.
+4 Experimental Evaluations
 In this chapter, we aim to verify the effectiveness of our proposed method in terms of performance preservation and memory reduction.
 Firstly, we introduce the experimental setup, including the selected model architectures, the baseline methods for comparison, and parameters such as the datasets used.
-Secondly, we verify the effects of our method on memory control and throughput.
+Secondly, we verify the performances of our method on memory control and throughput.
 Moreover, we compare the generation quality of different methods in the field of long texts and tasks.
@@
 4.1 Experiment Setup
-The backbone models selected in the experiment are some small-scale open-sourced LLMs, including LLaMA-2 7B, LLaMa3-8B-Instruct, and Mistral-7B-Instruct.
+The backbone models selected in the experiment are some small-scale open-sourced LLMs, including LLaMA2-7B, LLaMA3-8B-Instruct, and Mistral-7B-Instruct.
 The experimental equipment is an NVIDIA A800-80GB GPU.
@@
 The baselines for comparison are different KV cache management methods. Specifically, they include:
-  • Full KV (FKV): Caches all keys and values per token in each layer, serving as the vanilla baseline with maximal computational load but theoretically optimal performance.
+  • Full KV (FullKV): Caches all keys and values per token in each layer, serving as the vanilla baseline with maximal computational load but theoretically optimal performance.
   • Heavy Hitter Oracle (H2O): Dynamically manages KV cache using cumulative attention thresholds to balance recent tokens and Heavy Hitters (H2).
@@
 Implements attention rectification to compensate for distribution shifts caused by merged states, enabling sublinear cache growth while preserving contextual coherence.
-  • ZeroMerge (Ours): Builds upon merging paradigm but eliminates parametric dependencies through heuristic-based 0-shot merging.
+  • ZSMerge (Ours): Builds upon merging paradigm but eliminates parametric dependencies through heuristic-based zero-shot merging.
 Strictly enforces preset cache bounds via two-phase compression: 1) Attention sink preservation for initial/recent tokens, 2) Distance-aware similarity clustering for intermediate tokens.
 Minimizes embedding distortion through revised attention scoring that accounts for merged token positional densities, achieving better generalization across compression ratios than prior learning-based methods.
-4.2 Inference efficiency and resource consumption
+4.2 Inference Efficiency Gain
+To demonstrate the benefits of ZSMerge in improving inference efficiency, we conducted two proof-of-concept experiments.
+The first (Figure 3) compares the performance of full KV caching and ZSMerge under increasing sequence lengths in a specific inference case.
+The second (Table 1) evaluates ZSMerge against full KV caching and other baseline methods under various workloads, including different sequence lengths, batch sizes, and model sizes.
+Below, we summarize the memory and throughput improvements achieved by ZSMerge.
 Figure 3: VRAM Utilization and Decoding Throughput Across Sequence Lengths.
-ZeroMerge enforces constant memory footprint (43GB) and sustains 9 tokens/sec decoding rate beyond 54K tokens, eliminating OOM via dynamic KV cache compression.
-We evaluate the long - context processing efficiency on a single NVIDIA A800 GPU with 80GB VRAM (Video Random Access Memory).
-The experiment uses the LLaMA2 - 7B model, which has 32 layers, 32 attention heads, and an embedding dimension of 4096, in float32 precision.
-The benchmark is composed of a synthetic QA task, where the cache tokens range from 20K to 54K.
-For the baseline measurements, we use full KV caching without compression.
-In contrast, the ZeroMerge method enforces a token cache budget of 18K.
-The original Llama2-7B model (float32) required 25GB VRAM for parameter loading alone.
-During prefill phase, processing 20K prompt tokens generated 20GB KV cache across 32 layers, totally consuming 56% of total available VRAM (Fig.3).
-In baseline inference, KV cache grew linearly with sequence length at 1MB/token.
-By 54K tokens, cache memory dominated the mast of VRAM, triggering OOM errors.
-Concurrently, decoding throughput degraded from 9 tokens/sec (20K context) to 4 tokens/sec (52K context)—due to the increased computation of attention score.
-By enforcing a 18K token cache ceiling through two-phase compression (attention sink preservation + similarity-based merging), ZeroMerge: Reduced prefill-phase KV cache to 18GB (10% reduction); Maintained constant 43GB VRAM usage during decoding (vs.
-baseline’s 45 to 79 GB linear growth); Sustained stable throughput at 9 tokens/sec across 20K–54K token contexts; Eliminated OOM constraints through dynamic cache compression.
-4.3 Workload-Scalable KV Efficiency
-We evaluate ZeroMerge against three baseline methods (FullKV, H2O, LESS) on an NVIDIA A800 GPU (80GB VRAM) using LLaMA2-7B (float32) with diverse workloads: sequence lengths (2K–54K tokens), batch sizes (4–16), and document-QA tasks.
-Each method is tested under identical decoding constraints (temperature=0.0, greedy sampling) to isolate computational overhead.
-Metrics include throughput (tokens/sec), latency (sec), and OOM thresholds.
-Table 1: Workload-Scalable KV Cache Compression: ZeroMerge Outperforms Baselines in Throughput (tokens/sec) and Latency (seconds) Across Sequence Lengths and Batch Sizes.
+ZSMerge enforces constant memory footprint (43GB) and sustains 9 tokens/sec decoding rate beyond 54K tokens, eliminating out-of-memory (OOM) via dynamic KV cache compression.
+Table 1: Workload-Scalable KV Cache Compression: ZSMerge Outperforms Baselines in Throughput (tokens/sec) and Latency (seconds) Across Sequence Lengths and Batch Sizes.
 Seq.length Batch size Model size throughput(tokens/S) $\Uparrow$ / latency(S) $\Downarrow$
-FullKV H2O (5%) Less (5%) ZeroMerge (5%)
+FullKV H2O (5%) Less (5%) ZSMerge (5%)
 4096+4096 4 7B 62.4 / 262.1 43.9 / 372.5 15.0 / 1086.2 77.0 / 212.7
@@
 4096+4096 16 13B OOM - - 178.2 / 367.6
-As demonstrated in Table 1, ZeroMerge consistently outperforms existing methods across diverse workload configurations.
-At moderate batch sizes (batch=4, 4K+4K tokens), ZeroMerge achieves 23.4% higher throughput (77.0 vs 62.4 tokens/sec) and 18.9% lower latency (212.7s vs 262.1s) compared to FullKV, while H2O and LESS suffer 29.7% and 75.9% performance degradation respectively.
-Under high-throughput demands (batch=8, 4K+4K tokens), ZeroMerge successfully processes requests where both H2O and LESS encounter OOM errors, delivering 124% throughput improvement (146.7 vs 65.5 tokens/sec) over FullKV.
-Most notably, at large batch sizes (batch=16, 2K+2K tokens), ZeroMerge sustains 2.1× higher throughput (281.9 vs 133.1 tokens/sec) with 52.8% latency reduction (116.2s vs 246.2s), showcasing its unique capability to decouple computational load from sequence length.
-These results validate ZeroMerge’s dual advantage, eliminating memory bottlenecks through strict cache budgeting.
-4.4 Inference quality
+4.2.1 Memory Reduction
+Specific Case Analysis
+In the baseline setup, the LLaMA2-7B model required 25GB of VRAM for parameter loading, with an additional 20GB KV cache generated during the prefill phase for 20K tokens.
+This linear growth in KV cache size, at 1MB per token, led to OOM errors as sequence lengths approached 54K tokens.
+ZSMerge, constrained by an 18K token cache budget, reduced KV cache size by 10% during the prefill phase and maintained VRAM usage at a constant 43GB during decoding.
+This prevented the baseline’s linear memory growth (up to 79GB) and completely eliminated OOM errors, enabling efficient long-context processing.
+4.2.2 Throughput Improvement
+Micro-Level Analysis
+Decoding throughput for the baseline dropped from 9 tokens/sec at 20K tokens to 4 tokens/sec at 54K tokens due to increasing attention computation overhead.
+ZSMerge, in contrast, maintained a consistent throughput of 9 tokens/sec across the same sequence length range by dynamically merging less relevant tokens while preserving critical attention information.
+Macro-Level Analysis
+Under diverse workloads, ZSMerge consistently outperformed the baseline and other methods.
+For moderate batch sizes (batch=4), it achieved 23.4% higher throughput and 18.9% lower latency than full KV caching.
+At higher batch sizes (batch=8), ZSMerge handled requests that caused OOM errors for other methods, delivering a 124% throughput improvement.
+For large batch sizes (batch=16), it sustained 2.1× higher throughput and halved latency compared to full KV caching.
+In summary, ZSMerge effectively reduces memory usage and maintains high throughput across various workloads, outperforming full KV caching and other baselines.
+Its dynamic cache compression mechanism ensures scalability and efficient long-context processing without sacrificing performance.
+4.3 Inference Quality Loss
 We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics.
-The study compares three KV cache compression methods across two 7B models (Llama2, Falcon) at 20%, 10%, and 5% cache budgets.
+The experiment compares three KV cache compression methods across two 7B models (LLaMA2-7B and Falcon) at 20%, 10%, and 5% cache budgets.
 H2O balances recent tokens and heavy hitters per its original design, while LESS employs its recommended C4-trained merging network.
-ZeroMerge operates with $\alpha=0.6$ (influence ratio of residual cache ) and $\lambda=0.98$ (the decay rate of the historical contribution) without task-specific training.
-Our evaluation demonstrates that ZeroMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods.
-When compressing the KV cache to 5% of its original size, ZeroMerge maintains near-original performance on Llama2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions.
-This advantage becomes more pronounced at higher compression ratios, where ZeroMerge’s performance preservation significantly outperforms both baselines.
-On Falcon-7B, a model with distinct architectural characteristics, ZeroMerge demonstrates stronger generalization than parameter-dependent approaches.
-While LESS suffers severe quality degradation—likely due to its reliance on C4 dataset training—ZeroMerge retains over half of the baseline performance even at extreme 5% compression.
+For ZSMerge, the hyperparameters were set to static values based on empirical validation: the residual compensation intensity ratio ($\alpha$) was fixed at $0.6$ to balance performance across tasks, and the historical contribution decay rate ($\lambda$) was set to $0.98$ to ensure stable context propagation while prioritizing recent attention patterns.
+These values were selected through comparative analysis against other configurations (e.g., $\alpha\in[0.6,0.8]$ and $\lambda=0.98$ showed optimal trade-offs between compression efficiency and generation quality) and required no task-specific training.
+Our evaluation demonstrates that ZSMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods.
+When compressing the KV cache to 5% of its original size, ZSMerge maintains near-original performance on LLaMA2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions.
+This advantage becomes more pronounced at higher compression ratios, where ZSMerge’s performance preservation significantly outperforms both baselines.
+On Falcon-7B, a model with distinct architectural characteristics, ZSMerge demonstrates stronger generalization than parameter-dependent approaches.
+While LESS suffers severe quality degradation-likely due to its reliance on C4 dataset training-ZSMerge retains over half of the baseline performance even at extreme 5% compression.
 In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
-The results highlight two critical trends: First, higher compression ratios amplify ZeroMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies.
+The results highlight two critical trends: First, higher compression ratios amplify ZSMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies.
 Second, architecture-agnostic design enables consistent cross-model robustness, avoiding the domain adaptation challenges faced by data-driven methods like LESS.
 Table 2: Model-Agnostic Text Generation Quality under KV Cache Compression: ROUGE-1/2/L Scores for LLaMA2-7B and Falcon-7B Across 20%, 10%, and 5% Cache Budgets
-method Llama2 7B Falcon 7B
-FULL 30.59 / 11.34 / 25.50 27.06 / 8.79 / 22.39
+method LLaMA2-7B Falcon-7B
+Full KV 30.59 / 11.34 / 25.50 27.06 / 8.79 / 22.39
 20% H2O 30.83 / 11.43 / 25.71 24.18 / 7.47 / 20.14
 20% LESS 30.47 / 11.23 / 25.44 24.90 / 7.97 / 20.76
-20% ZeroMerge 31.62 / 12.40 / 26.42 25.19 / 8.69 / 21.34
+20% ZSMerge 31.62 / 12.40 / 26.42 25.19 / 8.69 / 21.34
 10% H2O 30.18 / 11.32 / 25.28 13.03 / 3.16 / 11.02
 10% LESS 30.74 / 11.22 / 25.58 8.99 / 2.55 / 7.50
-10% ZeroMerge 31.83 / 12.47 / 26.75 20.92 / 6.74 / 17.50
+10% ZSMerge 31.83 / 12.47 / 26.75 20.92 / 6.74 / 17.50
 5% H2O 28.92 / 10.81 / 24.35 12.02 / 2.08 / 10.36
 5% LESS 29.98 / 11.09 / 25.02 7.75 / 1.15 / 6.76
-5% ZeroMerge 30.60 / 11.67 / 25.72 15.04 / 3.29 / 12.73
-Acknowledgements
-Do not include acknowledgements in the initial version of the paper submitted for blind review.
-If a paper is accepted, the final camera-ready version can (and usually should) include acknowledgements.
-Such acknowledgements should be placed at the end of the section, in an unnumbered section that does not count towards the paper page limit.
-Typically, this will include thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
+5% ZSMerge 30.60 / 11.67 / 25.72 15.04 / 3.29 / 12.73
+5 Conclusion
+This study addresses the critical challenge of memory and computational inefficiency in long-context language models through ZSMerge, a dynamic KV cache compression framework.
+By introducing head-level memory allocation, residual merging with compensated attention scoring, and architecture-agnostic adaptation, the method achieves sublinear memory growth while preserving generation quality across diverse workloads.
+Empirical validation demonstrates that ZSMerge reduces VRAM consumption by 82% at 54K tokens,triples inference efficiency under a 5% compression ratio, and outperforms eviction-based baselines by 12-34% in text quality metrics—all without task-specific training.
+The framework advances practical deployment of long-context LLMs by fundamentally decoupling memory constraints from sequence length, enabling efficient processing of extended dialogues and documents while maintaining semantic coherence.
 Impact Statement
-Authors are required to include a statement of the potential broader impact of their work, including its ethical aspects and future societal consequences.
-This statement should be in an unnumbered section at the end of the paper (co-located with Acknowledgements – the two may appear in either order, but both must be before References), and does not count toward the paper page limit.
-In many cases, where the ethical impacts and expected societal implications are those that are well established when advancing the field of Machine Learning, substantial discussion is not required, and a simple statement such as the following will suffice:
-“This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.”
-The above statement can be used verbatim in such cases, but we encourage authors to think about whether there is content which does warrant further discussion, as this statement will be apparent if the paper is later flagged for ethics review.
+This paper presents work whose goal is to advance the fields of machine learning and sustainable artificial intelligence.
+There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.

### v3

@@
 The linear growth of key-value (KV) cache memory and quadratic computational complexity in attention mechanisms pose significant bottlenecks for large language models (LLMs) in long-context processing.
-While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often incur irreversible information loss or require costly parameter retraining.
-To this end, we propose ZSMerge, a dynamic KV cache compression framework designed for efficient cache management, featuring three key operations: (1) fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) a residual merging mechanism that preserves critical context through compensated attention scoring, and (3) a zero-shot adaptation mechanism compatible with diverse LLM architectures without requiring retraining.
-ZSMerge significantly enhances memory efficiency and inference speed with negligible performance degradation across LLMs.
-When applied to LLaMA2-7B, it demonstrates a 20:1 compression ratio for key-value cache retention (reducing memory footprint to 5% of baseline) while sustaining comparable generation quality, coupled with triple throughput gains at extreme 54k-token contexts that eliminate out-of-memory failures.
+While existing KV cache optimization methods address these challenges through token pruning or feature merging, they often incur irreversible information loss or require costly retraining.
+To this end, we propose ZSMerge, a dynamic KV cache compression framework for efficient cache management, featuring three key operations: (1) fine-grained memory allocation guided by multi-dimensional token importance metrics at head-level granularity, (2) a residual merging mechanism that preserves critical context through compensated attention scoring, and (3) a zero-shot adaptation mechanism compatible with diverse LLM architectures without retraining.
+ZSMerge significantly enhances memory efficiency and inference speed with negligible performance degradation.
+When applied to LLaMA2-7B, it achieves a 20:1 compression ratio for KV cache retention (reducing memory footprint to 5% of baseline) while sustaining comparable generation quality and tripling throughput at extreme 54k-token contexts, eliminating out-of-memory failures.
+These gains extend across diverse LLM families and long-context benchmarks, with ZSMerge consistently surpassing state-of-the-art cache optimization methods across tasks such as summarization, QA, reasoning, and code completion.
 The code is available at https://github.com/SusCom-Lab/ZSMerge.
-Machine Learning, ICML
 1 Introduction
-The advancement of the Transformer architecture has revolutionized sequence data processing, with Large Language Models (LLMs) being a prime illustration.
-LLMs have achieved remarkable progress, surpassing human capabilities in diverse applications.
-While multi-turn dialogue and long-context understanding are critical scenarios, handling extended sequences remains challenging due to the linear expansion of the Key-Value (KV) cache (which stores intermediate attention keys and values during generation to avoid redundant computations) in proportion to text length.
-For instance, deploying a LLaMA2-7B model will consume about 26GB of GPU memory; when the token length reaches 32K, the KV cache alone occupies 32G of memory, becoming the primary memory consumption.
-This, along with the quadratic computational complexity, significantly restricts LLM integration and performance, thus motivating recent research efforts toward optimizing KV cache utilization for enhanced inference efficiency.
+The advancement of the Transformer architecture has revolutionized sequence data processing, with Large Language Models (LLMs) being a prime illustration (vaswani2017attention, ).
+LLMs have achieved remarkable progress, surpassing human performance in diverse applications.
+While multi-turn dialogue and long-context understanding are critical scenarios (yi2024survey, ; gao2023self, ; li2023loogle, ; liu2024lost, ), handling extended sequences remains challenging due to the linear expansion of the Key-Value (KV) cache, which stores intermediate attention keys and values during generation to avoid redundant computations (kwon2023efficient, ).
+For instance, deploying a LLaMA2-7B (sarah2024llama, ) model will consume about 26GB of GPU memory; when the token length reaches 32K, the KV cache alone occupies 32GB of memory, becoming the primary memory consumption.
+This, along with the quadratic computational complexity, significantly restricts LLM integration and performance, thus motivating recent research efforts toward optimizing KV cache utilization for enhanced inference efficiency (zhang2024unifying, ; sun2024shadowkv, ; wang2024squeezeattention, ).
 To mitigate the large memory footprint of the KV cache, various optimization strategies have emerged.
-Some modify the core model architecture (Sarah et al., 2024; Ashkboos et al., 2024) or employ quantization techniques for lower precision representation (Dettmers et al., 2022; Xiao et al., 2023a).
+Some modify the core model architecture (sarah2024llama, ; ashkboos2024slicegpt, ) or employ quantization techniques for lower precision representation (dettmers2022gpt3, ; xiao2023smoothquant, ).
 Other methods directly target the KV cache bottleneck during inference using context-aware techniques.
-Common among these is sparse approximation or token eviction, retaining only crucial tokens while discarding others (Zhang et al., 2024b, a); however, this can disrupt KV embeddings and cause information loss.
-To mitigate this, alternative context-aware techniques merge or fuse information from similar tokens (Dong et al., 2024), often using low-rank compression, though these may introduce architectural overhead or require model retraining.
+Common among these is sparse approximation or token eviction, retaining only crucial tokens while discarding others (zhang2024h2o, ; zhang2024pyramidkv, ); however, this can disrupt KV embeddings and cause information loss.
+To mitigate this, alternative context-aware techniques merge or fuse information from similar tokens (dong_get_2024, ), often using low-rank compression, though these may introduce architectural overhead or require model retraining.
 Ideally, an effective and practical KV cache management strategy should 1) strictly control the memory consumption to break the context length limits, 2) minimize the generation quality degradation, and 3) readily applicable across diverse LLM backbones without necessitating fine-tuning or architectural changes (zero-shot compatibility).
 Motivated by these goals, we propose ZSMerge, an efficient, comprehensive, and dynamic zero-shot KV management algorithm, with the following key features:
-  • Fine-grained memory allocation.
-ZSMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and its own data distribution characteristics to conduct fine-grained, head-level memory management.
+  • Fine-grained memory allocation: ZSMerge makes full use of the historical contribution of tokens, spatio-temporal characteristics, and intrinsic data distribution characteristics to conduct fine-grained, head-level memory management.
 This adaptability allows it to maintain superior generation quality compared to the sparse strategies and achieve performance comparable to retraining-dependent approaches, even when operating with significantly reduced memory budgets.
 This ensures effectiveness across diverse long-context tasks.
-  • Zero-shot and low-cost Integration.
-Designed for ease of use, ZSMerge does not introduce any model parameters and thus requires no retraining or model fine-tuning to adapt to different compression ratios or various task types.
-This also makes it compatible to various mainstream LLM architectures such as LLaMA, Falcon, and Mistral, thereby minimizing deployment overhead.
-  • Efficient and throughput-boosting strategy.
-Characterized by linear operational complexity, ZSMerge’s compression mechanism is computationally lightweight.
+  • Zero-shot and low-cost Integration: Designed for ease of use, ZSMerge does not introduce any model parameters and thus requires no retraining or model fine-tuning to adapt to different compression ratios or various task types.
+This also makes it compatible to various mainstream LLM architectures such as LLaMA, Falcon (almazrouei2023falcon, ), Mistral (jiang2023mistral, ), Qwen (yang2024qwen2, ), and Yi (aiYiOpenFoundation2025, ), thereby minimizing deployment overhead.
+  • Efficient throughput-boosting strategy: Characterized by linear operational complexity, ZSMerge’s compression mechanism is computationally lightweight.
 This efficiency brings significant throughput improvement, e.g., it helps achieve over triple the inference throughput compared to the original model at a token length of 54K.
@@
 2 Related Work
-Figure 1: Comparative analysis of KV cache optimization strategies: (A) Sparse approximation directly addresses memory-compute growth by evicting low-contribution tokens, albeit inducing attention distribution drift.
-(B) Token merge mitigates sparse caching errors, often by learning residuals implicitly or explicitly, yet introduces overhead from auxiliary networks or fine-tuning.
-(C) ZSMerge (proposed) achieves zero-shot compression competitive with methods requiring learned components, doing so without added parameters while demonstrating superior generalization.
-The escalating memory and computational demands of large language models (LLMs), particularly from KV cache growth in long-context scenarios, have spurred diverse optimization strategies.
-These are broadly categorized into context-agnostic and contextually adaptive approaches.
+KV cache growth in long-context LLMs creates severe memory and computational bottlenecks, motivating optimization strategies across two categories: context-agnostic and contextually adaptive approaches.
 2.1 Context-Agnostic Optimization
-Context-agnostic methods universally reduce resource consumption by altering the model’s structure or numerical representation, irrespective of the input sequence.
-Foundational techniques include Structural Compression via knowledge distillation (e.g., MiniLLM (Gu et al., 2024)) or matrix factorization (e.g., SliceGPT (Ashkboos et al., 2024)).
-Attention Mechanism Optimization, such as head pruning (Voita et al., 2019) and key-value sharing in Multi-Query (MQA (Shazeer, 2019)) and Grouped-Query Attention (GQA (Ainslie et al., 2023)), significantly cuts memory bandwidth.
-Simultaneously, Numerical Optimization tackles precision redundancy through techniques like post-training quantization (e.g., GPTQ (Frantar et al., 2022)), hardware-aligned quantization (e.g., INT-FlashAttention (Chen et al., 2024)), and exploiting low-rank properties (e.g., Linformer (Wang et al., 2020)).
-These strategies collectively alleviate memory bottlenecks via architectural simplification and precision reduction.
+Context-agnostic methods reduce resource consumption through structural or numerical modifications independent of input sequences.
+Key approaches include: (1) Structural compression via knowledge distillation (gu2024minillm, ) or matrix factorization (ashkboos2024slicegpt, ); (2) Attention optimization through head pruning (voita2019analyzing, ) and key-value sharing in Multi-Query (shazeer2019fast, ) and Grouped-Query Attention (ainslie2023gqa, ); (3) Numerical optimization via post-training quantization (frantar2022gptq, ), hardware-aligned quantization (chen2024int, ), and low-rank approximation (wang2020linformer, ).
 2.2 Contextually Adaptive Optimization
-While context-agnostic methods offer baseline efficiency, recent research increasingly aims to directly target the KV cache—a critical bottleneck in long-sequence tasks driven by storage growth and quadratic computational complexity.
-The observed sparsity in attention values (Ji et al., 2021; Deng et al., 2024) motivates dynamic, context-aware strategies.
-Figure 1 illustrates the main categories of these adaptive approaches:
-(A) Sparse Approximation (Eviction): This class directly addresses memory-compute growth by evicting low-contribution tokens, based on the premise that not all history is equally vital.
-Work like (Xiao et al., 2023b) identified ”attention sinks,” leading to methods like StreamingLLM (Xiao et al., 2023b) which retains initial tokens alongside a sliding window.
-LM-infinite (Han et al., 2024) uses specialized masks to balance local and global context.
-H2O (Zhang et al., 2024b) employs attention score thresholds for dynamic eviction, while PyramidKV (Zhang et al., 2024a) uses layer-wise attention patterns for progressive caching.
-Dynamic routing (Roy et al., 2021) also fits here.
-However, eviction strategies inevitably induce attention distribution drift and error propagation by permanently discarding tokens (Dao et al., 2022; Xiao et al., 2023b).
-(B) Token Merging: To mitigate information loss from eviction, token merging fuses similar tokens instead of discarding them.
-Pioneered in vision (e.g., ToMe (Bolya et al., 2022)), it applies to text as well.
-Dynamic Memory Compression adaptively merges spatially proximate tokens, often requiring fine-tuning.
-LESS (Dong et al., 2024) uses auxiliary networks to compress tokens with minimal attention impact.
+Context-aware methods directly target KV cache bottlenecks by exploiting attention sparsity (ji2021distribution, ; deng2024attention, ). Figure 1 illustrates three main approaches:
+Figure 1: Comparative analysis of contextually adaptive KV cache optimization strategies.
+(A) Sparse Approximation (Eviction): Low-contribution tokens are permanently discarded, reducing memory but risking error propagation.
+(B) Token Merging: Similar tokens are fused to mitigate information loss, but may require fine-tuning or auxiliary networks.
+(C) ZSMerge (Proposed): Achieves zero-shot "sparse+residual" compression without added parameters or tuning, balancing efficiency and performance.
+Sparse Approximation (Eviction): This class of methods directly tackles the growth of memory and computation by evicting low-contribution tokens, under the assumption that not all historical context is equally important.
+For example, StreamingLLM xiao2023efficient identifies "attention sinks" and retains both the initial tokens and a sliding context window.
+LM-infinite han2024lm uses specialized masks to balance local and global context.
+SnapKV li2024snapkv performs one-time KV cache pruning during the prefilling stage based on relatively consistent attention patterns, but lacks dynamic adjustment during decoding.
+H2O zhang2024h2o employs attention score thresholds for dynamic eviction, while PyramidKV zhang2024pyramidkv uses layer-wise attention patterns for progressive caching.
+Recent advances include FlexPrefill laiFlexPrefillContextAwareSparse2025 , which applies context-aware sparse attention specifically during prefilling.
+Dynamic routing roy2021efficient also fits here.
+However, eviction strategies inevitably induce attention distribution drift and error propagation by permanently discarding tokens dao2022flashattention ; xiao2023efficient .
+Token Merging: To mitigate information loss from eviction, token merging fuses similar tokens instead of discarding them.
+Pioneered in vision (e.g., ToMe bolya2022token ), it applies to text as well.
+Dynamic Memory Compression nawrot2024dynamic adaptively merges spatially proximate tokens, often requiring fine-tuning.
+LESS dong_get_2024 uses auxiliary networks to compress tokens with minimal attention impact.
+ThinK xuThinKThinnerKey2024 takes a different approach by applying query-driven pruning at the channel level, orthogonal to token-level compression.
 While merging mitigates sparse caching errors by learning residuals (implicitly or explicitly), it often introduces overhead from auxiliary networks or fine-tuning, complicating deployment.
-(C) ZSMerge (Proposed): Addressing the limitations of both prior approaches, we propose ZSMerge.
-It achieves zero-shot compression competitive with sparse+residual methods without added parameters, while demonstrating superior generalization across task domains and compression ratios.
-ZSMerge offers a parameter-free, tuning-free approach, balancing efficiency and generation performance.
+Training-Dependent and System-Level Approaches: Recent work has explored complementary directions to token-level compression.
+Training-dependent methods include NSA (Native Sparse Attention) yuanNativeSparseAttention2025 , which introduces MLP-based compressor modules trained jointly with the attention mechanism, and MoBA (Mixture of Block Attention) luMoBAMixtureBlock2025 , which performs block-level pruning via gating with auxiliary routing networks.
+System-level approaches like OmniKV haoOmniKVDynamicContext2024 focus on dynamic context selection and offloading strategies for efficient long-context processing.
+InfiniPot kimInfiniPotInfiniteContext2024 adopts periodic block-wise compression with a fixed KV cache budget to reduce runtime overhead.
+While powerful, these methods require retraining, architectural modifications, or specialized system configurations.
+ZSMerge (Proposed): Addressing the limitations of prior approaches, we propose ZSMerge.
+It achieves zero-shot compression competitive using the "sparse+residual" method without added parameters, while demonstrating superior generalization across task domains and compression ratios.
+ZSMerge offers a parameter-free, tuning-free approach that is complementary to training-dependent and system-level methods, balancing efficiency and generation performance.
 3 ZSMerge Methodology
@@
 Consider an $L$-layer transformer with multi-head attention mechanisms. For a target attention head at decoding step $T$, the cached Key and Value matrices are defined as:
-$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top}\in\mathbb{R}^{T\times d},$ (1)
-$\quad\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top}\in\mathbb{R}^{T\times d},$
+$\mathbf{K}_{T}=[\mathbf{k}_{1},\mathbf{k}_{2},\ldots,\mathbf{k}_{T}]^{\top}\in\mathbb{R}^{T\times d},~\mathbf{V}_{T}=[\mathbf{v}_{1},\mathbf{v}_{2},\ldots,\mathbf{v}_{T}]^{\top}\in\mathbb{R}^{T\times d},$ (1)
 where $\mathbf{k}_{t},\mathbf{v}_{t}\in\mathbb{R}^{d}$ represent the Key/Value vectors for the $t$-th token.
 For the query vector $\mathbf{q}_{T}\in\mathbb{R}^{d}$ at position $T$, the scaled dot-product attention computes output $\mathbf{o}^{(T)}$ via:
-$a_{t}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{s=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d})},$ (2)
-$\quad\mathbf{o}^{(T)}$ $=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}$
+$a_{t}^{(T)}=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}{\sum_{i=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})},~\mathbf{o}^{(T)}=\sum_{t=1}^{T}a_{t}^{(T)}\mathbf{v}_{t}.$ (2)
 This formulation establishes the baseline for analyzing cache compression effects on attention distribution fidelity.
@@
 $B=B_{p}+B_{c}+B_{r},$ (3)
-where $B_{p}$, $B_{c}$, and $B_{r}$ govern context preservation, residual absorption, and proximity maintenance, respectively. The compressed cache matrices aggregate three complementary components:
+where $B_{p}$, $B_{c}$, and $B_{r}$ govern proximity maintenance, context preservation, and residual absorption, respectively.
+This tripartite division enables a nuanced approach to compression.
+Our empirical findings (see Appendix C.2) indicate that allocating the largest share of the budget to $B_{c}$ and $B_{p}$ is generally effective, as it prioritizes the retention of core information and local context.
+In contrast, $B_{r}$ is typically assigned a smaller budget to handle fine-grained residual adjustments.
+The compressed cache matrices therefore integrate three complementary components:
 $\mathbf{K}_{B}=[\mathbf{K}_{p}\|\mathbf{K}_{c}\|\mathbf{K}_{r}],\quad\mathbf{V}_{B}=[\mathbf{V}_{p}\|\mathbf{V}_{c}\|\mathbf{V}_{r}],$ (4)
-[…]
+with $\|$ denoting row-wise concatenation.
 The proximity component $(\mathbf{K}_{p},\mathbf{V}_{p})$ preserves the latest $B_{p}$ tokens, capturing local attention patterns.
 The context component $(\mathbf{K}_{c},\mathbf{V}_{c})$ retains top-$B_{c}$ tokens ranked by contribution scores $\mathbf{s}^{(T)}\in\mathbb{R}^{T}$, which quantify contextual saliency.
-The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through learned transformations.
-This configuration subsumes eviction-based methods when $B_{r}=0$, while enabling continuous interpolation between sparse and dense representations through budget reallocation.
+The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ dynamically merges $B_{r}$ historically evicted tokens through attention-weighted aggregation transformations.
+The residual component $(\mathbf{K}_{r},\mathbf{V}_{r})$ maintains $B_{r}$ dedicated token slots, representing a compressed summary of previously evicted tokens.
+During generation, the residual cache is updated dynamically, progressively merging newly expelled tokens into the existing compressed representation.
+This configuration subsumes eviction-based methods when $B_{r}=0$.
 3.3 Contribution Evaluation
@@
 $s_{t}^{(T)}=\begin{cases}\lambda s_{t}^{(T-1)}+a_{t}^{(T)},&T>0\\ 0,&\text{otherwise}\end{cases},$ (5)
-where the decay factor $\lambda\in[0,1]$ operates as a temporal discounting parameter analogous to reinforcement learning credit assignment, governing the exponential decay of historical attention contributions.
-We recommend setting it to a relatively large value to balance the contributions of long-term and short-term attention scores.
-Throughout this paper, $\lambda$ is configured as $0.98$ unless otherwise specified, as scores from 300 steps earlier become attenuated to approximately 0.002 through exponential decay ($0.98^{300}\approx 0.002$).
+where the decay factor $\lambda\in[0,1]$ acts as a temporal discounting parameter analogous to reinforcement learning credit assignment, controlling the exponential decay of historical attention contributions.
+In practice, we find that the performance is not highly sensitive to the precise value of $\lambda$.
+We therefore fix $\lambda=0.98$ throughout this paper for simplicity, which yields a smooth balance between long-term and short-term contributions.
 3.4 Residual Token Merging
@@
   2. 2.
-    Update the selected slot through momentum-based fusion:
+    Update the selected slot via incremental mean aggregation:
     $\mathbf{k}_{\hat{r}}\leftarrow\frac{w_{\hat{r}}\mathbf{k}_{\hat{r}}+\mathbf{k}_{t}}{w_{\hat{r}}+1},\quad\mathbf{v}_{\hat{r}}\leftarrow\frac{w_{\hat{r}}\mathbf{v}_{\hat{r}}+\mathbf{v}_{t}}{w_{\hat{r}}+1}$ (7)
@@
 Figure 2 illustrates the dynamic cache evolution under budget parameters $B_{r}=2$, $B_{c}=4$, and $B_{p}=3$ during sequence length expansion from $T=12$ to $T=17$.
-Figure 2: KV cache operation process with $B_{r}=2$, $B_{c}=4$, $B_{p}=3$, and sequence length expansion $T=12\rightarrow 17$: residual merging (2 slots), context preservation (top-4) and proximity maintenance (latest 3).
+3.5 Attention Output Stabilization
 Algorithm 1 ZSMerge Online Compression
@@
 14:    else
-15:              Merge token $(\mathbf{k}_{\hat{c}},\mathbf{v}_{\hat{c}})$ into $\mathbf{K}_{r}$ via (6)-(7)
+[…]
 16:    end if
@@
 19:end for
-3.5 Attention Output Stabilization
+Figure 2: KV cache operation process
 Algorithm 1 progressively merges low-saliency tokens through dynamic cache updates.
-To faithfully reconstruct the original attention distribution from compressed representations, we introduce a compensated attention scoring mechanism.
+To more accurately reconstruct the attention distribution from compressed representations, we introduce a compensated attention scoring mechanism.
 The revised attention computation evolves from Eq.
 2 to:
-$\hat{a}_{t}^{(T)}$ $=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{i=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d}+\alpha\log w_{i}\right)},$ (8)
-$\mathbf{\hat{o}}^{(T)}$ $=\sum_{t=1}^{T}\hat{a}_{t}^{(T)}\mathbf{v}_{t},$
+$\hat{a}_{t}^{(T)}=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{i=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d}+\alpha\log w_{i}\right)},~\mathbf{\hat{o}}^{(T)}=\sum_{t=1}^{T}\hat{a}_{t}^{(T)}\mathbf{v}_{t},$ (8)
 where $w_{i}$ represents the fusion count of token $i$ (with $w_{i}=1$ for uncompressed tokens).
-The hyperparameter $\alpha\in(0,1]$ regulates residual compensation intensity, with empirical analysis suggesting an optimal range of $\alpha\in[0.6,0.8]$ for balanced performance.
-Notably, setting $\alpha=0$ reduces the merging algorithm to a degenerate case equivalent to a pure eviction policy, eliminating residual compensation entirely.
+The hyperparameter $\alpha\in[0,1]$ provides a soft transition, with $\alpha=0$ degenerating to an eviction-like policy.
+Its concrete influence is demonstrated in the appendix, and we fix $\alpha=1$ throughout our experiments.
 With Eq. 8, we address two key challenges:
@@
   • Attention Mass Conservation: The compensation term preserves the relative attention mass between compressed and uncompressed tokens, ensuring residual compensation does not suppress critical uncompressed components.
-Theorem 3.1.
+Theorem 1.
 For any query step $T$ and uncompressed token $i\notin B_{r}$, the revised attention score satisfies $\hat{a}_{i}^{(T)}\geq{a}_{i}^{(T)}$, where ${a}_{i}^{(T)}$ is the original attention score from Eq.
 2.
-Proof.
-Let $r\in B_{r}$ be a residual slot merging $\{r_{1},...,r_{w_{r}}\}$ tokens. For compressed tokens, we establish an upper bound on their attention numerator:
-$\text{num}(\hat{a}_{r}^{(T)})$ $=\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}+\alpha\log w_{r}\right)$ (9)
-$\leq w_{r}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}\right)$
-$=w_{r}\exp\left(\frac{1}{w_{r}}\sum_{m=1}^{w_{r}}\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right)$
-$\leq\sum_{m=1}^{w_{r}}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right),$
-where the first inequality uses $\alpha\leq 1$, and the second applies Jensen’s inequality to the convex exponential function. For uncompressed tokens ($w_{i}=1$), we derive:
-$\hat{a}_{i}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t\notin B_{r}}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})+\sum_{r\in B_{r}}\text{num}(\hat{a}_{r}^{(T)})}$ (10)
-$\geq\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}={a}_{i}^{(T)}.$
-as the denominator contains compressed tokens’ upper-bounded contributions. ∎
+The proof is provided in Appendix A.
 This theoretical guarantee ensures that uncompressed tokens retain their relative dominance in attention allocation despite cache compression, even as the denominator accounts for upper-bounded contributions from compressed tokens.
@@
 4 Experimental Evaluations
-In this chapter, we aim to verify the effectiveness of our proposed method in terms of performance preservation and memory reduction.
-Firstly, we introduce the experimental setup, including the selected model architectures, the baseline methods for comparison, and parameters such as the datasets used.
-Secondly, we verify the performances of our method on memory control and throughput.
-Moreover, we compare the generation quality of different methods in the field of long texts and tasks.
-Finally, we discuss the trade-off between memory, time, and performance as well as the impact of relevant hyperparameters on the model.
+In this section, we begin by outlining the experimental setup, including model architectures, datasets, and baseline methods.
+We then assess efficiency improvements in terms of memory and speed.
+Next, we provide a detailed numerical error analysis to highlight the role of representation bias correction.
+Finally, we evaluate generation quality under diverse cache budgets, across multiple task types, and over different base model series, while comparing against a broad set of baselines.
 4.1 Experiment Setup
-The backbone models selected in the experiment are some small-scale open-sourced LLMs, including LLaMA2-7B, LLaMA3-8B-Instruct, and Mistral-7B-Instruct.
-The experimental equipment is an NVIDIA A800-80GB GPU.
-For the dataset, we choose LongBench, which is a comprehensive benchmark dataset specifically designed for evaluating the performance of different methods on tasks that involve long-context inputs.
-It contains a wide variety of natural language processing tasks with different lengths of text sequences, covering multiple domains such as news, novels, and academic papers, thus providing a rich and diverse set of test cases for our experiments.
-The baselines for comparison are different KV cache management methods. Specifically, they include:
-  • Full KV (FullKV): Caches all keys and values per token in each layer, serving as the vanilla baseline with maximal computational load but theoretically optimal performance.
-  • Heavy Hitter Oracle (H2O): Dynamically manages KV cache using cumulative attention thresholds to balance recent tokens and Heavy Hitters (H2).
-Leverages layer-wise attention score averaging across queries to identify critical tokens, achieving adaptive compression while maintaining task-specific semantics.
-  • LESS: Introduces recurrent KV state synthesis through layer-aware dynamic merging.
-Combines temporal locality exploitation (retaining recent tokens) with spatial compression via parameterized merging operators.
-Implements attention rectification to compensate for distribution shifts caused by merged states, enabling sublinear cache growth while preserving contextual coherence.
-  • ZSMerge (Ours): Builds upon merging paradigm but eliminates parametric dependencies through heuristic-based zero-shot merging.
-Strictly enforces preset cache bounds via two-phase compression: 1) Attention sink preservation for initial/recent tokens, 2) Distance-aware similarity clustering for intermediate tokens.
-Minimizes embedding distortion through revised attention scoring that accounts for merged token positional densities, achieving better generalization across compression ratios than prior learning-based methods.
+Model Coverage: Our evaluation encompasses diverse modern LLM architectures for broad compatibility.
+Core experiments use LLaMA2-7B, Falcon-7B, and Mistral-7B-Instruct on NVIDIA A800-80GB GPU.
+We extend evaluation to LLaMA-3.1-8B-Instruct, Qwen2.5-7B-Instruct, and Yi-6B (Appendix C.4), spanning different attention mechanisms including Multi-Query Attention (MQA).
+Benchmark Diversity: We use synthetic datasets for efficiency evaluation and multiple public benchmarks for quality assessment: LongBench (bai2024longbench, ) (21 tasks, 6 categories), XSum (narayanDontGiveMe2018, ) (summarization), InfiniteBench (zhang-etal-2024-bench, ) (100K+ tokens), and GSM-Infinite-8k (zhouGSMInfiniteHowYour2025, ) (mathematical reasoning).
+Extended results are in Appendix C.
+Baselines: We compare ZSMerge against representative KV cache management methods:
+  • FullKV: Stores all key-value pairs at every layer, providing the vanilla baseline with maximal memory and compute overhead, but serving as the upper bound for task performance.
+  • StreamingLLM (Stream) (xiao2023efficient, ): Retains both early prompt tokens and a fixed sliding window of recent tokens.
+By designating "attention sinks," it ensures semantic anchors remain available.
+This static rule is efficient and robust but lacks adaptivity to task-specific token importance.
+  • SnapKV (li2024snapkv, ): Performs a one-time cache pruning during prefilling, leveraging early attention distributions to discard less relevant tokens.
+It eliminates runtime overhead and achieves efficiency, but cannot adapt if token importance shifts during generation.
+  • H2O (zhang2024h2o, ): Maintains adaptivity via cumulative attention thresholds that dynamically retain heavy hitters alongside recency-biased tokens.
+Layer-wise averaging of attention scores captures persistent importance, enabling balanced compression while preserving key semantics.
+  • LESS (dong_get_2024, ): Introduces dynamic KV state synthesis through recurrent merging.
+It combines recency preservation with similarity-based compression, then applies attention rectification to correct distortions introduced by merging.
+This allows sublinear cache growth without severe loss of contextual coherence.
+Extended Baseline Coverage: Beyond the core baselines, our comprehensive evaluation includes additional state-of-the-art methods to provide thorough comparative analysis.
+These include OmniKV (haoOmniKVDynamicContext2024, ) for attention-guided compression, InfLLM (xiaoInfLLMTrainingFreeLongContext2024, ) for infinite-length modeling, Minference (jiangMInference10Accelerating2024, ) for efficient attention approximation, and FlexPrefill (laiFlexPrefillContextAwareSparse2025, ) for flexible prefilling strategies.
+Detailed comparisons with these advanced baselines on challenging benchmarks such as InfiniteBench and GSM-Infinite-8k are presented in Appendix C.4, demonstrating ZSMerge’s competitive performance against the latest compression techniques.
 4.2 Inference Efficiency Gain
@@
 FullKV H2O (5%) Less (5%) ZSMerge (5%)
-4096+4096 4 7B 62.4 / 262.1 43.9 / 372.5 15.0 / 1086.2 77.0 / 212.7
-4096+4096 8 7B 65.5 / 500.3 OOM OOM 146.7 / 223.2
-2048+2048 16 7B 133.1 / 246.2 86.1 / 380.1 OOM 281.9 / 178.36
-4096+4096 16 13B OOM - - 178.2 / 367.6
+1024+1024 8 7B 177.8 / 46.1 104.9 / 78.1 48.8 / 167.9 161.6 / 50.7
+2048+2048 8 7B 110.8 / 147.9 72.2 / 227.0 25.1 / 654.2 163.2 / 100.4
+2048+2048 16 7B 133.1 / 246.2 86.1 / 380.1 OOM 281.9 / 178.4
+4096+4096 8 13B OOM OOM OOM 110.8 / 295.7
+2048+2048 2 13B 43.0 / 95.2 25.5 / 160.4 12.7 / 322.7 31.3 / 131.0
+2048+2048 4 13B 59.7 / 137.2 40.7 / 201.2 16.5 / 497.5 61.5 / 133.3
+4096+4096 4 13B 37.1 / 441.8 24.9 / 657.9 OOM 60.0 / 273.2
+4096+4096 16 13B OOM OOM OOM 178.2 / 367.6
 4.2.1 Memory Reduction
@@
 4.2.2 Throughput Improvement
-Micro-Level Analysis
 Decoding throughput for the baseline dropped from 9 tokens/sec at 20K tokens to 4 tokens/sec at 54K tokens due to increasing attention computation overhead.
 ZSMerge, in contrast, maintained a consistent throughput of 9 tokens/sec across the same sequence length range by dynamically merging less relevant tokens while preserving critical attention information.
-Macro-Level Analysis
-Under diverse workloads, ZSMerge consistently outperformed the baseline and other methods.
-For moderate batch sizes (batch=4), it achieved 23.4% higher throughput and 18.9% lower latency than full KV caching.
-At higher batch sizes (batch=8), ZSMerge handled requests that caused OOM errors for other methods, delivering a 124% throughput improvement.
-For large batch sizes (batch=16), it sustained 2.1× higher throughput and halved latency compared to full KV caching.
-In summary, ZSMerge effectively reduces memory usage and maintains high throughput across various workloads, outperforming full KV caching and other baselines.
-Its dynamic cache compression mechanism ensures scalability and efficient long-context processing without sacrificing performance.
+ZSMerge consistently outperformed baselines across diverse workloads, achieving superior throughput while avoiding OOM errors that plagued other methods.
+Notably, for memory-intensive scenarios (e.g., 13B model, 4096+4096 seq, batch 16), ZSMerge was the only method capable of processing requests (178.2 tokens/s), demonstrating its scalability advantage through dynamic compression.
+4.3 Numerical Error Analysis
+To further substantiate the role of representation bias correction, we conduct a numerical error analysis that isolates the effect of different compression strategies.
+In particular, we measure the relative error of attention outputs under varying cache budget constraints, comparing pure eviction-based compression against our residual merging approach.
+This setup directly reflects the practical conditions under which compression-induced representation shifts occur.
+Figure 4: Numerical error analysis comparing eviction vs.
+residual merging across compression ratios.
+The log-scale visualization demonstrates that ZSMerge’s residual merging consistently reduces approximation error relative to pure eviction, with benefits amplified under aggressive compression ($\leq 20\%$ cache size).
+As shown in Figure 4, under aggressive compression scenarios ($\leq 20\%$ cache size), ZSMerge’s residual merging demonstrates substantial error reduction: 60.5% at 20% cache size, 43.8% at 10%, and 37.4% at 5%.
+The most dramatic improvement occurs at 50% compression, where residual merging achieves 89.1% error reduction ($2.72\times 10^{-4}$ vs.
+$2.50\times 10^{-3}$).
+This empirical validation confirms that the residual compensation mechanism effectively preserves attention distribution coherence, preventing the error propagation characteristic of pure eviction strategies.
+The results show that residual merging consistently yields lower approximation error than eviction-based methods, especially under tight memory budgets where aggressive compression is required.
+This analysis empirically validates the representation-bias-correction hypothesis underlying ZSMerge.
+Additional validation on the efficiency of residual compensation and its connection to Eq.
+8 is provided in Appendix C.2.
 4.4 Inference Quality Loss
-We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics.
-The experiment compares three KV cache compression methods across two 7B models (LLaMA2-7B and Falcon) at 20%, 10%, and 5% cache budgets.
-H2O balances recent tokens and heavy hitters per its original design, while LESS employs its recommended C4-trained merging network.
-For ZSMerge, the hyperparameters were set to static values based on empirical validation: the residual compensation intensity ratio ($\alpha$) was fixed at $0.6$ to balance performance across tasks, and the historical contribution decay rate ($\lambda$) was set to $0.98$ to ensure stable context propagation while prioritizing recent attention patterns.
-These values were selected through comparative analysis against other configurations (e.g., $\alpha\in[0.6,0.8]$ and $\lambda=0.98$ showed optimal trade-offs between compression efficiency and generation quality) and required no task-specific training.
-Our evaluation demonstrates that ZSMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods.
-When compressing the KV cache to 5% of its original size, ZSMerge maintains near-original performance on LLaMA2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions.
-This advantage becomes more pronounced at higher compression ratios, where ZSMerge’s performance preservation significantly outperforms both baselines.
-On Falcon-7B, a model with distinct architectural characteristics, ZSMerge demonstrates stronger generalization than parameter-dependent approaches.
-While LESS suffers severe quality degradation-likely due to its reliance on C4 dataset training-ZSMerge retains over half of the baseline performance even at extreme 5% compression.
-In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
-The results highlight two critical trends: First, higher compression ratios amplify ZSMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies.
-Second, architecture-agnostic design enables consistent cross-model robustness, avoiding the domain adaptation challenges faced by data-driven methods like LESS.
-Table 2: Model-Agnostic Text Generation Quality under KV Cache Compression: ROUGE-1/2/L Scores for LLaMA2-7B and Falcon-7B Across 20%, 10%, and 5% Cache Budgets
+4.4.1 Impact of Cache Budget
+Table 2: Text Generation Quality under KV Cache Compression: ROUGE-1/2/L Scores for LLaMA2-7B and Falcon-7B Across 20%, 10%, and 5% Cache Budgets
 method LLaMA2-7B Falcon-7B
@@
 5% ZSMerge 30.60 / 11.67 / 25.72 15.04 / 3.29 / 12.73
+We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics.
+Experiments compare three KV-cache compression methods across two 7B-parameter models (LLaMA2-7B and Falcon-7B) under 20%, 10%, and 5% cache budgets.
+H2O follows its original design, balancing recent tokens and heavy hitters, while LESS uses its recommended merging network trained on the C4 dataset raffel2020exploring .
+For ZSMerge, budget allocation matches LESS ($B_{r}=2$, with $B_{p}$ and $B_{c}$ sharing the remaining budget), and other hyperparameters are fixed as previously described.
+Our evaluation demonstrates that ZSMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods.
+As shown in Table 2, when compressing the KV cache to 5% of its original size, ZSMerge maintains near-original performance on LLaMA2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions.
+This advantage becomes more pronounced at higher compression ratios, where ZSMerge’s performance preservation significantly outperforms both baselines.
+On Falcon-7B, a model with distinct architectural characteristics, ZSMerge demonstrates stronger generalization than parameter-dependent approaches.
+While LESS suffers severe quality degradation-likely due to its reliance on C4 dataset training.
+ZSMerge retains over half of the baseline performance even at extreme 5% compression.
+In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
+The results highlight two critical trends: First, higher compression ratios amplify ZSMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies.
+Second, its training-free nature enables consistent robustness across different model architectures, avoiding the domain adaptation challenges faced by data-driven methods such as LESS that require training data to tune auxiliary modules.
+4.4.2 Generalization across Task Types
+Table 3: Results on LongBench benchmark including code completion (CODE), few-shot learning(FSHOT), multi-document QA (MDQA), single-document QA (SDQA), summarization (SUMM), and synthetic reasoning (SYNC) tasks.
+Method
+CODE
+FSHOT
+MDQA
+SDQA
+SUMM
+SYNC
+LlaMA2-7B
+FullKV 65.19 52.32 11.48 17.50 15.15 5.01
+Cache size=512
+H2O 19.63 7.37 4.48 4.05 3.51 2.16
+Stream 61.26 47.63 8.35 11.60 7.07 4.32
+SnapKV 63.06 51.51 10.33 15.85 12.28 5.50
+ZSMerge 63.02 51.42 10.36 15.67 12.17 5.34
+Cache size=1024
+H2O 29.12 20.34 5.92 7.47 6.70 2.97
+Stream 62.94 49.98 8.16 11.80 7.18 4.45
+SnapKV 64.45 52.02 11.30 16.65 13.30 5.09
+ZSMerge 64.41 52.06 11.37 16.63 13.23 5.02
+Mistra-7B
+FullKV 62.10 63.09 37.34 38.94 26.14 66.67
+Cache size=512
+Stream 57.90 53.68 27.04 23.92 19.80 32.17
+SnapKV 60.46 62.24 34.75 36.86 22.18 64.83
+ZSMerge 60.48 62.17 34.62 36.83 22.09 64.50
+Cache size=1024
+Stream 60.13 56.01 27.80 25.10 21.25 34.17
+SnapKV 61.60 62.36 35.49 37.30 23.73 66.50
+ZSMerge 61.59 62.42 35.43 37.31 23.64 66.50
+We evaluate the effectiveness of ZSMerge on the LongBench benchmark, which covers a diverse set of task types, including code completion, few-shot learning, multi-document QA, single-document QA, summarization, and synthetic reasoning tasks.
+Experiments are conducted on two backbone models: LLaMA2-7B and Mistral-7B, with cache size constraints set to 512 and 1024.
+For the Mistral-7B model, the H2O baseline encountered out-of-memory (OOM) errors and could not be included in the comparison.
+More details about the experimental setup and validation results on specific datasets are provided in Appendix C.
+Overall, the results in Table 3 show that ZSMerge consistently matches or closely tracks the performance of FullKV across most task types, despite operating under strict cache constraints.
+Compared to Stream and H2O, ZSMerge demonstrates substantially higher accuracy, especially in complex and memory-intensive tasks like multi-document QA and summarization.
+Interestingly, while SnapKV benefits from a one-time pruning strategy during the prefilling phase which allows it to retain a larger cache during decoding, ZSMerge still achieves comparable or even slightly better performance in several tasks, thanks to its adaptive two-phase compression and positional-aware merging strategy.
+These findings highlight ZSMerge’s strong generalization and robustness across heterogeneous workloads, making it a promising solution for real-world deployment in memory-constrained scenarios.
+Additional results with richer benchmarks and broader baseline comparisons are provided in Appendix C.4.
 5 Conclusion
-This study addresses the critical challenge of memory and computational inefficiency in long-context language models through ZSMerge, a dynamic KV cache compression framework.
-By introducing head-level memory allocation, residual merging with compensated attention scoring, and architecture-agnostic adaptation, the method achieves sublinear memory growth while preserving generation quality across diverse workloads.
-Empirical validation demonstrates that ZSMerge reduces VRAM consumption by 82% at 54K tokens,triples inference efficiency under a 5% compression ratio, and outperforms eviction-based baselines by 12-34% in text quality metrics—all without task-specific training.
-The framework advances practical deployment of long-context LLMs by fundamentally decoupling memory constraints from sequence length, enabling efficient processing of extended dialogues and documents while maintaining semantic coherence.
-Impact Statement
-This paper presents work whose goal is to advance the fields of machine learning and sustainable artificial intelligence.
-There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+This study tackles the memory and computational inefficiency of long-context language models with ZSMerge, a dynamic KV cache compression framework.
+Through head-level memory allocation, residual merging with compensated attention scoring, and architecture-agnostic adaptation, ZSMerge achieves sublinear memory growth while preserving generation quality.
+Empirical results show an 82% VRAM reduction at 54K tokens, a threefold increase in inference efficiency under a 5% compression ratio, and 12–34% higher text quality metrics over eviction-based baselines, all without task-specific training.
+Additionally, by enabling efficient deployment on resource-constrained devices, ZSMerge reduces energy consumption and carbon emissions, advancing the sustainability of LLM applications in real-world scenarios.
+Appendix A Proof of Theorem 1
+Proof of Theorem 1.
+Let $r\in B_{r}$ be a residual slot merging $\{r_{1},...,r_{w_{r}}\}$ tokens. For compressed tokens, we establish an upper bound on their attention numerator:
+$\text{num}(\hat{a}_{r}^{(T)})$ $=\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}+\alpha\log w_{r}\right)$ (9)
+$\leq w_{r}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r}/\sqrt{d}\right)$
+$=w_{r}\exp\left(\frac{1}{w_{r}}\sum_{m=1}^{w_{r}}\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right)$
+$\leq\sum_{m=1}^{w_{r}}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{r_{m}}/\sqrt{d}\right),$
+where the first inequality uses $\alpha\leq 1$, and the second applies Jensen’s inequality [50] to the convex exponential function. For uncompressed tokens ($w_{i}=1$), we derive:
+$\hat{a}_{i}^{(T)}$ $=\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t\notin B_{r}}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})+\sum_{r\in B_{r}}\text{num}(\hat{a}_{r}^{(T)})}$ (10)
+$\geq\frac{\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d})}{\sum_{t=1}^{T}\exp(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d})}={a}_{i}^{(T)}.$
+as the denominator contains compressed tokens’ upper-bounded contributions. ∎
+Appendix B Implementation Details
+This section presents the implementation details of ZSMerge.
+The KV cache compression framework is built upon the Transformers library.
+To minimize deviations from the original framework and reduce redevelopment complexity, only the forward propagation function was replaced globally.
+As a result, a single process cannot simultaneously hold two instances with different compression modes.
+However, the compression mode can be easily switched without creating a new instance by calling the change_modechange\_mode method.
+Our framework currently supports replacing the scaled_dot_product_attentionscaled\_dot\_product\_attention function for the LLaMA, Falcon, and Mistral model families, as this operation is widely used across various inference scenarios.
+The initialization of the attention score $s$ based on the full history of attention scores during the prefilling stage imposes a substantial computational burden.
+In certain long-sequence tasks (such as the LongBench experiments), we introduce a hyperparameter, window_sizewindow\_size, to limit the range of timesteps considered during the initialization of $s$, following the approach used in SnapKV.
+This optimization has a minimal impact on generation quality but significantly accelerates the prefilling process.
+Appendix C Extended Experiments
+C.1 Latency and Throughput Evaluation Across Sequence Lengths and Batch Sizes
+Table 4 presents additional experimental results on workload-scalable KV cache compression.
+The validation of the LESS and H2O frameworks was conducted using the code provided at https://github.com/hdong920/LESS.
+The results were obtained from a single experiment run, as the observed conclusions were clear and consistent.
+In short-sequence and low-batch-size settings, the FullKV method shows slight advantages, as compression introduces additional computational overhead.
+However, in other scenarios, ZSMerge demonstrates significant performance gains, even when compared to other compression methods.
+Table 4: Workload-Scalable KV Cache Compression: ZSMerge Outperforms Baselines in Throughput (tokens/sec) and Latency (seconds) Across Sequence Lengths and Batch Sizes.
+Seq.length Batch size Model size throughput(tokens/S) $\Uparrow$ / latency(S) $\Downarrow$
+FullKV H2O (5%) Less (5%) ZSMerge (5%)
+1024+1024 4 7B 117.5 / 34.9 83.5 / 49.0 22.7 / 180.6 81.5 / 50.3
+1024+1024 8 7B 177.8 / 46.1 104.9 / 78.1 48.8 / 167.9 161.6 / 50.7
+2048+2048 8 7B 110.8 / 147.9 72.2 / 227.0 25.1 / 654.2 163.2 / 100.4
+2048+2048 16 7B 133.1 / 246.2 86.1 / 380.1 OOM 281.9 / 178.4
+4096+4096 4 7B 62.4 / 262.1 43.9 / 372.5 15.0 / 1086.2 77.0 / 212.7
+4096+4096 8 7B 65.5 / 500.3 OOM OOM 146.7 / 223.2
+4096+4096 16 7B OOM OOM OOM 271.6 / 241.3
+8192+4096 4 7B 38.9 / 420.8 OOM OOM 74.3 / 220.4
+8192+8192 4 7B 33.1 / 989.3 OOM OOM 78.3 / 418.5
+8192+4096 8 7B OOM OOM OOM 132.0 / 248.2
+8192+8192 8 7B OOM OOM OOM 142.6 / 459.6
+256+256 2 13B 62.5 / 8.2 49.2 / 10.4 31.1 / 16.5 30.7 / 16.7
+512+512 2 13B 61.4 / 16.7 43.5 / 23.6 27.1 / 37.8 31.1 / 32.9
+1024+1024 2 13B 54.4 / 37.7 33.2 / 61.6 16.8 / 121.8 31.1 / 65.8
+2048+2048 2 13B 43.0 / 95.2 25.5 / 160.4 12.7 / 322.7 31.3 / 131.0
+2048+2048 4 13B 59.7 / 137.2 40.7 / 201.2 16.5 / 497.5 61.5 / 133.3
+4096+4096 4 13B 37.1 / 441.8 24.9 / 657.9 OOM 60.0 / 273.2
+4096+4096 8 13B OOM OOM OOM 110.8 / 295.7
+4096+4096 16 13B OOM OOM OOM 178.2 / 367.6
+4096+8192 16 13B OOM OOM OOM 666.3 / 196.7
+4096+4096 32 13B OOM OOM OOM 397.5 / 329.7
+8192+8192 4 13B OOM OOM OOM 617.0 / 53.1
+8192+8192 8 13B OOM OOM OOM 642.3 / 102.0
+8192+4096 16 13B OOM OOM OOM 466.7 / 140.4
+8192+8192 16 13B OOM OOM OOM 812.6 / 161.3
+8192+8192 32 13B OOM OOM OOM OOM
+C.2 Hyperparameter Sensitivity Validation
+As a training-free framework for KV cache compression, our method introduces several hyperparameters to enhance flexibility and provide a smooth transition to classical sparsity-based methods.
+We conduct comprehensive sensitivity analysis to offer empirical recommendations for practical deployment scenarios.
+Budget Distribution Strategy: Our framework follows a hierarchical budget allocation strategy.
+First, the proximity maintenance budget $B_{p}$ is controlled by the cache tail ratio ($B_{p}/B$).
+Then, a small portion of the remaining budget is allocated to the residual budget $B_{r}$, controlled by the cache dense parameter ($B_{r}/(B-B_{p})$).
+Finally, the remaining budget is distributed to the context preservation budget $B_{c}$.
+Experimental Setup: We conduct sensitivity analysis on LLaMA2-7B using the XSUM summarization task.
+We fix an anchor configuration and systematically vary each hyperparameter to isolate its individual impact on performance, measured by ROUGE-1, ROUGE-2, and ROUGE-L scores.
+  • Proximity Maintenance Ratio ($B_{p}/B$): This parameter governs the allocation between proximity maintenance and context preservation.
+Our analysis reveals that extreme partitions significantly hinder performance.
+Values below 0.3 or above 0.7 show notable degradation, with ROUGE-1 scores dropping by 4-8% at the extremes (0.1 and 0.9).
+The optimal range lies between 0.3 and 0.7, with peak performance around 0.5, consistent with established methods like H2O.
+We recommend setting this ratio to 0.5 for balanced performance.
+  • Residual Budget Ratio ($B_{r}/(B-B_{p})$): The residual budget demonstrates the effectiveness of our token merging operation.
+When $B_{r}=0$, the method degrades to a pure eviction strategy, showing 6-11% performance drops across all ROUGE metrics.
+Small positive values (0.01-0.03) provide stable benefits, with 0.02 showing optimal performance.
+Higher values (0.05-0.08) compress the context budget excessively, leading to performance degradation of 7-18%.
+We recommend setting this ratio to 0.02 to achieve stable benefits while preserving sufficient context budget.
+  • Scale Factor ($\alpha$): This parameter controls the influence of merged cache tokens.
+Our analysis shows that increasing the scale factor from 0.0 to 1.0 progressively improves performance, with ROUGE scores improving by 1-5%.
+Setting $\alpha=0$ degenerates the method to standard sparse approaches, while $\alpha=1.0$ provides optimal performance.
+This validates the effectiveness of our token merging operation.
+We restrict the value to 1.0 in our standard experiments.
+Figure 5: Hyperparameter Sensitivity Analysis: (top) Proximity Maintenance Ratio ($B_{p}/B$), (middle) Residual Budget Ratio ($B_{r}/(B-B_{p})$), (bottom) Scale Factor ($\alpha$)
+Key Findings: Figure 5 illustrates the sensitivity patterns across all three hyperparameters.
+The analysis confirms that ZSMerge demonstrates robust performance across most hyperparameter configurations.
+The method shows particular sensitivity to extreme budget allocations but maintains stable performance within recommended ranges.
+The effectiveness of token merging is clearly demonstrated through the consistent improvements observed when enabling residual budgets and scale factors, distinguishing our approach from pure eviction-based methods.
+Practical Recommendations: For deployment scenarios, we recommend the following configuration: $B_{p}/B=0.5$, $B_{r}/(B-B_{p})=0.02$, and $\alpha=1.0$.
+This configuration provides robust performance while maintaining compatibility with existing sparse attention frameworks.
+C.3 Details on LongBench benchmark
+Dataset: We evaluate ZSMerge on the LongBench benchmark, a comprehensive suite for assessing long-context understanding in LLMs. The benchmark comprises 21 tasks spanning six categories:
+  • Single-document QA
+  • Multi-document QA
+  • Summarization
+  • Few-shot learning
+  • Synthetic tasks
+  • Code completion
+The datasets (English and Chinese) feature context lengths of 5,000-15,000 tokens and are standardized for automated evaluation [43].
+Evaluation Framework: Our experiments utilize the benchmarking methods implemented in the KVCache-Factory repository (https://github.com/Zefan-Cai/KVCache-Factory).
+This framework supports various KV cache compression methods, including PyramidKV, SnapKV, StreamingLLM, and H2O.
+It is compatible with attention mechanisms such as Flash Attention v2 and SDPA, allowing for efficient evaluation under different memory constraints .
+Models and Configuration: We conduct experiments on two backbone models: LLaMA2-7B and Mistral-7B.
+To simulate memory-constrained scenarios, we set the cache size constraints to 512 and 1024 tokens.
+Notably, for the Mistral-7B model, the H2O baseline encountered out-of-memory (OOM) errors and was excluded from the comparison.
+Table 5: Performance Comparison of KV Cache Compression Methods on LongBench Tasks
+Method
+2wikimqa
+dureader
+gov_report
+hotpotqa
+lcc
+lsht
+multi_news
+multifieldqa_en
+multifieldqa_zh
+musique
+narrativeqa
+passage_count
+passage_retrieval_en
+passage_retrieval_zh
+qasper
+qmsum
+repobench-p
+samsum
+trec
+triviaqa
+vcsum
+LlaMA2-7B
+FullKV 10.54 23.35 27.16 7.77 68.13 20.25 3.01 23.93 18.78 4.26 17.33 1.50 5.52 8.00 9.94 20.55 62.25 32.11 68.00 88.92 9.89
+Cache size=512
+H2O 7.33 3.72 1.18 4.81 22.06 0.00 1.79 5.62 1.90 2.05 5.53 2.36 4.03 0.08 3.16 10.99 17.19 3.08 18.50 7.90 0.10
+Stream 9.43 13.97 0.98 6.60 64.73 16.67 0.29 17.21 11.82 3.41 11.45 1.88 5.12 5.96 5.93 18.74 57.79 30.34 56.00 87.51 8.28
+SnapKV 10.70 17.87 18.18 8.33 66.35 17.25 2.61 21.76 17.00 4.41 17.39 2.75 6.26 7.50 7.26 19.98 59.76 33.96 67.50 87.34 8.37
+ZSMerge 10.72 17.98 17.62 8.34 66.32 17.25 2.65 21.52 16.59 4.41 17.30 2.75 6.26 7.00 7.29 20.05 59.71 33.62 67.50 87.29 8.37
+Cache size=1024
+H2O 8.49 6.41 4.25 6.08 38.60 0.50 2.91 11.14 4.84 2.70 6.34 2.23 6.67 0.00 7.55 19.51 19.65 30.86 19.50 30.52 0.14
+Stream 9.29 13.33 0.99 6.49 66.61 17.00 0.91 16.99 12.10 3.51 11.47 1.62 3.92 7.81 6.66 19.05 59.27 31.87 62.50 88.54 7.77
+SnapKV 10.67 21.78 22.54 8.02 67.64 18.75 2.74 22.56 18.04 4.73 17.49 2.54 5.98 6.75 8.52 20.58 61.26 32.97 68.00 88.38 7.33
+ZSMerge 10.67 22.12 22.05 7.97 67.63 18.75 2.82 22.39 18.07 4.73 17.55 2.54 5.76 6.75 8.49 20.67 61.19 33.13 68.00 88.38 7.39
+Mistral-7B-Instruct-v0.3
+FullKV 39.01 32.38 34.89 49.37 61.56 40.25 27.83 52.88 32.26 28.58 29.07 5.50 98.00 96.50 41.58 25.77 62.63 47.51 76.00 88.59 16.08
+Cache size=512
+Stream 31.86 17.64 22.10 41.05 59.37 18.50 23.20 29.91 15.60 17.60 24.21 6.00 81.00 9.50 25.95 20.25 56.42 43.79 65.50 86.95 13.65
+SnapKV 38.72 24.02 25.85 49.53 60.32 37.75 24.96 54.05 28.27 26.72 28.79 5.00 96.00 93.50 36.34 24.08 60.60 46.75 75.00 89.44 13.82
+ZSMerge 38.56 23.68 25.47 49.67 60.21 37.75 24.85 53.98 28.38 26.58 29.01 5.00 95.00 93.50 35.94 24.22 60.76 46.67 75.00 89.28 13.81
+Cache size=1024
+Stream 32.65 17.17 24.59 43.35 61.04 21.25 25.48 31.16 16.46 18.03 24.81 5.50 82.50 14.50 27.95 20.81 59.21 45.59 68.50 88.71 14.11
+SnapKV 38.86 26.13 28.30 49.14 61.34 38.25 26.72 52.64 29.73 27.81 29.09 5.50 98.00 96.00 37.76 25.13 61.86 46.22 76.00 88.99 14.76
+ZSMerge 38.86 25.89 28.05 49.14 61.35 38.25 26.67 52.64 29.50 27.81 29.09 5.50 98.00 96.00 38.03 25.00 61.83 46.42 76.00 88.99 14.82
+Llama-3.1-8B-Instruct
+FullKV 16.39 31.45 34.13 15.93 65.06 40.50 26.70 27.02 20.03 9.97 21.06 7.34 72.79 76.30 12.80 22.44 57.46 43.56 70.00 91.37 16.14
+Cache size=512
+SnapKV 15.26 23.65 25.02 16.17 63.93 40.00 24.19 25.97 20.43 8.71 20.00 7.72 72.28 79.51 11.10 22.94 54.27 42.34 67.50 91.67 13.79
+ZSMerge 14.83 23.56 25.24 16.36 64.45 40.00 24.21 24.98 19.24 8.97 21.73 7.35 71.25 74.57 10.45 22.70 55.08 43.27 66.00 91.47 14.23
+Cache size=1024
+SnapKV 15.02 25.37 27.55 16.54 64.41 40.00 25.49 27.63 19.94 9.73 20.21 6.24 72.40 74.99 11.16 23.31 55.97 42.96 69.50 91.39 14.10
+ZSMerge 15.01 25.13 27.82 16.35 64.33 40.00 25.65 26.27 19.63 8.88 20.49 7.25 72.29 76.07 11.12 22.88 57.17 43.52 69.00 91.08 14.77
+The experimental results (Table 5) reveal several key patterns.
+Across both LLaMA2-7B and Mistral-7B models, the uncompressed FullKV baseline achieves the highest performance but serves primarily as an upper-bound reference rather than a practical solution.
+When comparing compression methods under memory constraints, ZSMerge and SnapKV demonstrate superior capability in preserving model accuracy compared to StreamingLLM and H2O, particularly in challenging scenarios with cache sizes limited to 512 or 1024 tokens.
+The Mistral-7B model consistently outperforms LLaMA2-7B, showing particularly strong results in question answering and retrieval tasks, where it achieves near-perfect scores in some cases.
+This performance gap highlights Mistral’s architectural advantages for long-context processing.
+Meanwhile, LLaMA2 struggles more noticeably with certain synthetic tasks and Chinese language datasets, suggesting limitations in its multilingual and reasoning capabilities.
+Cache size plays a measurable but not decisive role in performance.
+While increasing the cache from 512 to 1024 tokens provides modest improvements, ZSMerge maintains competitive accuracy even with the smaller cache, demonstrating its efficiency.
+In contrast, H2O shows severe degradation under constrained settings, failing completely on some tasks.
+Task-specific analysis indicates that summarization and code-related tasks benefit most from methods that better preserve context, like ZSMerge and SnapKV.
+Retrieval-focused tasks, however, show less sensitivity to cache size, with Mistral achieving consistently high scores regardless of compression.
+Overall, ZSMerge emerges as a balanced solution, delivering near-FullKV performance while operating efficiently within strict memory limits.
+To further evaluate the applicability of ZSMerge under realistic deployment settings, we benchmarked it using the LLaMA-3.1-8B-Instruct model on the LongBench suite.
+As shown in Table 5, we report results exclusively for our method, since other contextually adaptive baselines (e.g., StreamingLLM, H2O) currently do not provide support for LLaMA-3 architecture.
+Despite this, ZSMerge exhibits robust performance under constrained cache settings (512 and 1024 tokens), achieving scores close to the uncompressed FullKV baseline across a wide range of tasks.
+In particular, ZSMerge preserves strong performance on summarization and retrieval tasks, indicating that our token merging strategy can mitigate MQA’s context sensitivity without auxiliary modules or task-specific tuning.
+These results reinforce the generalization and plug-and-play capability of ZSMerge, even when deployed with newer model architectures featuring aggressive attention simplification.
+C.4 Extended Architecture Validation
+To validate our zero-shot compatibility claims, we extended ZSMerge evaluation to modern LLM architectures.
+This section presents comprehensive results on LLaMA3 and discusses ongoing experiments with additional contemporary models.
+InfiniteBench Long-Context Evaluation
+We successfully implemented ZSMerge on LLaMA3-8B architecture, demonstrating that our framework maintains compatibility with stronger, more recent backbones.
+The implementation required no architectural modifications or hyperparameter retuning, confirming the architecture-agnostic design principles of our approach.
+To address concerns about evaluation scope, we conducted comprehensive experiments on InfiniteBench [45] using LLaMA3-8B, a demanding benchmark featuring contexts exceeding 100K tokens.
+This evaluation provides critical validation under extreme long-context scenarios that stress-test compression methods beyond conventional limits.
+Table 6: InfiniteBench Results: ZSMerge vs. State-of-the-Art Methods on LLaMA3-8B
+Method En.Dia En.MC En.Sum Math.Find Retrieve.KV Retrieve.Number Retrieve.PassKey Zh.QA Avg
+FullKV 37.37 23.70 21.72 55.55 73.73 98.00 100.00 25.47 54.69
+H2O 43.17 23.33 21.30 63.30 26.36 72.52 98.40 24.63 46.63
+InfLLM 24.67 15.85 16.91 50.85 0.00 98.00 100.00 34.87 42.64
+OmniKV 30.83 23.33 21.99 55.00 48.67 98.00 100.00 24.83 50.33
+Streaming LLM 14.63 13.88 20.31 40.10 0.00 5.83 2.73 17.67 14.39
+Minference 24.95 21.55 20.91 62.11 17.03 75.84 57.27 22.65 37.79
+FlexPrefill 33.56 23.47 21.44 70.51 53.53 98.17 99.49 27.99 53.52
+ZSMerge 36.44 22.79 21.34 55.16 67.96 98.00 100.00 24.84 52.95
+The InfiniteBench results (Table 6) demonstrate several critical findings:
+Performance Preservation:
+ZSMerge achieves 96.8% of FullKV performance (52.95 vs.
+54.69 average), maintaining quality despite compression.
+This validation occurs with context lengths of at least 100K tokens that test compression robustness.
+Competitive Positioning:
+The method outperforms traditional eviction-based approaches (H2O: 46.63, InfLLM: 42.64) and matches the performance of recent specialized methods (FlexPrefill: 53.52, OmniKV: 50.33).
+ZSMerge performs well in retrieval tasks (Retrieve.KV: 67.96), demonstrating information preservation for memory-intensive operations.
+Task-Specific Analysis:
+The method shows strength in:
+  • Retrieval tasks: Scores of 100.00 on PassKey and 98.00 on Number retrieval
+  • Long-form reasoning: Performance on English Dialogue (36.44) and Math Finding (55.16)
+  • Cross-lingual capability: Performance on Chinese QA (24.84), indicating multilingual robustness
+Extended Baseline Comparison with SnapKV
+We conducted evaluations with SnapKV on models that were not originally supported, including Qwen2.5 [19], Yi-1.5 [20], and LLaMA-3.1.
+Implementation Adaptations:
+For comparison, we implemented technical adaptations including GQA model support for Grouped Query Attention models (LLaMA-3 and Qwen2.5) by averaging grouped query scores to align them with KV heads, Qwen compatibility by resolving interface differences in the RoPE method to maintain compatibility with the original Qwen behavior, and architecture extension by enabling SnapKV support for Qwen2.5-7B, Yi-1.5-6B, and LLaMA-3.1-8B-Instruct models.
+Experimental Setup:
+We evaluated on the GSM-Infinite-8k [46] benchmark across three difficulty levels (symbolic, medium, hard).
+The methodology included length bucketing where test samples were bucketed by input length to account for tokenizer variations despite the nominal 8k limit, adaptive cache budgets using cache budgets of 9k, 4k, and 2k tokens for samples with input lengths >10k, 5k–10k, and <5k tokens respectively, and ensuring all methods used identical experimental conditions and evaluation metrics.
+The results are presented in Table 7.
+Table 7: Extended Baseline Comparison: ZSMerge vs. SnapKV on GSM-Infinite-8k
+Model Method Symbolic Medium Hard
+Qwen2.5-7B-Instruct FullKV 9.35% 17.12% 12.04%
+SnapKV 0.00% 3.60% 8.33%
+ZSMerge 4.66% 9.91% 9.26%
+Yi-1.5-6B FullKV 0.00% 5.83% 7.46%
+SnapKV 0.00% 4.85% 6.47%
+ZSMerge 0.00% 5.34% 5.97%
+LLaMA-3.1-8B-Instruct FullKV 20.24% 11.65% 13.93%
+SnapKV 16.74% 10.19% 12.93%
+ZSMerge 15.38% 11.17% 13.43%
+Key Findings:
+The results demonstrate that ZSMerge consistently matches or outperforms SnapKV across all tested models and difficulty levels:
+  • Qwen2.5-7B: ZSMerge shows superior performance on symbolic and medium difficulty tasks, with competitive results on hard tasks.
+  • Yi-1.5-6B: ZSMerge achieves slightly better performance across medium and hard difficulties while maintaining identical symbolic task performance.
+  • LLaMA-3.1-8B: ZSMerge demonstrates consistent advantages across all difficulty levels, particularly excelling in symbolic reasoning tasks.
+These results validate that ZSMerge maintains its effectiveness when compared against state-of-the-art baselines across diverse modern architectures, confirming the robustness and generalizability of our approach.